A Complete Bibliography of Journal of the Royal Statistical Society, Series D (The Statistician):

1960–1969

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

27 July 2020 Version 1.02

Title word cross-reference

3 × 3 [370]. X̄ [143]. F [165]. n [153].

-Dimensions [153].

1 [168, 191]. 17th [411]. 1949 [499]. 1958 [173]. 1966 [466].

360 [484].

A.I.S. [46]. A.I.S./G.P.O. [46]. Absence [42]. Abstracts [579]. Abuse [268]. Academic [350]. Acceptance [43]. Accidents [113]. Achievement [180].

Achievements [110]. Ackoff [168, 215]. Activities [274]. Adaptive [474]. Address [137, 56]. Adequacy [49]. Adjustment [475]. Adkins [371]. Adler [252]. Administration [255]. Administrative [89]. Advanced [534, 114, 593]. Advancement [33]. Advertising [180, 5, 181, 182, 179]. Affairs [408]. Affecting [291]. Against [150]. Agricultural [266]. Agriculture [263, 187]. Aid [279, 317]. Aids [444, 312]. Air [364, 367]. Aircraft [570]. Aitcheson [578]. Aitken [150]. Alan [356]. Alexander [116]. Algebra [404, 300]. Alison [173, 499, 596]. Allied [321]. Among [264]. Analysing [83]. Analysis [223, 282, 579, 597, 31, 593, 254, 475, 572, 265, 311, 309, 452, 476, 125, 354, 47, 432, 232, 589, 463, 399, 127, 126, 290, 123, 231, 227, 373, 588, 247, 14, 162, 220, 124, 559, 599, 511, 395, 561, 216, 337, 50, 384, 514]. Analysis-Ideas [220]. Analysts [208]. Analytical [381]. Anatomy [588]. Anderson [535]. Andree [518]. Announcements [100, 118, 132, 155, 177, 193, 205, 221, 241, 259, 269, 304, 341, 359, 376, 389, 409, 427]. Annoyance [570]. Annual [411, 410, 466, 605, 19, 65]. Anson [439]. Answer [5]. appear [563]. Application [309, 107, 462]. Applications [381, 404, 352, 90, 115, 388, 99, 34, 69, 438, 336]. Applied [494, 415, 575, 458, 603, 322, 514]. Applying [277]. Appointments [223, 282]. Appreciation [144]. Approach [180, 105, 371, 220]. Approaches [303]. Arctic [379]. Arising [74]. Art [295]. Arthur [152]. Ascher [484]. Aspects [494, 181]. Assessment [379, 475, 293, 276, 465]. Association [572, 33, 10]. Astronomy [190]. Asymptotic [321]. Atta [250]. Attempt [5]. Attitude [64, 67]. Austin [202]. Autocodes [108]. Autoinstructional [294]. Automatic [434, 332]. Available [290]. Aylmer [144].

B [403, 297, 354, 115, 217, 442, 188]. Back [11, 27, 39, 53, 70, 84, 95, 120, 134, 157, 178, 194, 206, 222, 242, 260, 270, 284, 305, 326, 343, 361, 377, 390, 412, 429, 453, 471, 489, 505, 526, 542, 552, 568, 584, 608]. Bailey [352]. Barlow [421]. Barnett [321]. Barrie [600]. Bartholomew [537]. Bartlett [99]. Barton [148]. Based [310]. Basic [256, 302, 199]. Basis [142]. Bayesian [315, 595, 594]. Beckenbach [322]. Beer [532]. Behaviour [303]. Behavioural [371]. Bell [49]. Belson [49]. Bergstrom [601]. Berkeley [190]. Bernard [593]. Beta [9]. Between [165]. Beveridge [183]. Beyer [483]. Bharucha [69]. Bharucha-Reid [69]. Bias [368]. Bibliographic [293]. Bibliography [173, 499, 596, 256, 176, 402]. Binomial [165]. Biological [76, 404, 78, 6, 423]. Biology [190]. Biomedical [302]. BIOMETRIKA [8]. Biostatistics [577]. Bird [378]. Birds [380]. Blending [392]. Blom [9]. Board [57, 276]. Book [51, 302, 320, 321, 322, 323, 324, 388, 403, 420, 498, 534, 579, 594, 597, 423, 8, 22, 23, 26, 35, 49, 301, 371, 481, 338, 173, 499, 596, 187, 439, 513, 593, 130, 117, 131, 168, 190, 191, 240, 250, 254, 258, 297, 316, 374, 482, 497, 517, 535, 537, 560, 563, 577, 591, 38, 440, 515, 602, 82, 293, 355, 372, 267, 153, 317, 354, 385, 404, 422, 464, 536, 561, 592, 36, 68, 444, 401, 81, 115, 147, 78, 79, 99, 169, 170, 192, 237, 313, 396, 397, 398, 519, 520, 600, 601, 80, 575]. Book [576, 441, 484, 33, 66, 216, 256, 319, 337, 382, 483, 303, 564, 463, 148, 149, 150, 174, 175, 176, 201, 202, 217, 218, 238, 239, 257, 294, 295, 315, 352, 353, 386, 387, 399, 442, 443, 518, 268, 9, 25, 34, 69, 299, 114, 172, 24, 152, 402, 296, 219, 598, 83, 116, 154, 171, 199, 204, 255, 113, 98, 188, 189, 314, 595, 50, 151, 356, 373, 400, 421, 200, 438, 52, 339, 485, 298, 516, 335, 336, 210, 215, 252, 318, 383, 384, 514, 37, 67, 419, 565, 604, 203, 292, 48, 220, 251, 465]. Book [253, 300, 462, 603, 559, 562, 578, 599]. Box [107, 106, 417, 109]. BP [456]. Bradford [169, 202]. Breeding [379]. Britain [33]. British [33, 66, 378, 267, 298]. Bronchitis [365]. Brown [419]. Brownlee [117]. Bruce [595]. Bryan [219]. Bryant [50]. Buckland [256, 293]. Budin [255]. Building [576]. Bulmer [382]. Business [320, 22, 301, 82, 36, 204, 384, 420, 313, 80].

C

[423, 49, 439, 250, 482, 577, 267, 317, 519, 576, 150, 294, 171, 50, 151, 421, 438, 318, 383, 559]. Calculation [368]. Calculations [108, 338, 24]. Campling [598]. Can [598]. Capture [558]. Case [150]. Categorical [589]. Census [73]. Centre [456]. Chains [594]. Chakravarti [575]. Chambers [33]. Chance [148]. Change [528, 33, 419]. Changes [3]. Changing [74]. Characteristic [293, 336, 38]. Chemical [88]. Chemistry [381]. Chernoff [37]. Chi [209]. Chiang [577]. Chief [41]. Child [416]. Choice [511]. Choices [597]. Christ [482]. Chronic [365]. Classes [368, 546, 138]. Classification [434]. Classroom [373]. Clinical [202]. Co [598]. Coal [57, 62, 58]. Cochran [560]. Coefficient [64]. Coinage [33]. Collected [535, 113]. College [548]. Colliery [60, 59]. Colman [294]. Combinatorial [322, 564, 148]. Comment [141]. Commerce [33]. Commercial [89, 90]. Commodities [332]. Company [106, 484, 138, 109]. Comparison [477, 479, 195]. Comparisons [291, 251]. Complex [511]. Component [309, 127]. Components [480]. Computational [279, 480]. Computer [481, 279, 369, 484, 74, 395, 294, 518]. Computerized [561]. Computers [92, 478, 246]. Computing [587]. Concentration [60, 59, 66]. Concepts [250, 383]. Conduct [530, 509]. Conference [563, 46]. Congestion [131]. Connor [335]. Considerations [530]. Consignments [418]. Construction [207, 601, 557, 571]. Consumers [163]. Contingency [370]. Continuous [392]. Contributions [217]. Control [328, 579, 143, 106, 108, 330, 288, 573, 105, 107, 277, 109, 329, 299, 356]. Cook [492]. Correspondence [325, 145, 547, 406, 446, 574, 197, 211, 166, 128, 424, 521, 96, 111, 280, 233, 351, 185]. Corrigenda [234]. Corrugated [276]. Cotton [225]. Council [410, 466, 605, 556]. Count [30]. Course [498, 153, 68, 115]. Courses [140]. Coutie [318]. Cox [147, 399, 443, 35, 257]. Cramer [26, 516].

Crime [75]. Critical [579]. Crowe [388]. Cuming [439]. Cumulative [319]. Curves [321, 317, 290]. Cybernetic [239].

D [320, 321, 35, 371, 338, 187, 130, 297, 537, 385, 147, 78, 66, 148, 257, 295, 315, 399, 443, 188, 336, 318]. Daniel [403]. Data [558, 367, 511, 395, 480, 83, 373]. David [148, 189, 251]. Davies [318]. Debate [350]. Decimal [33]. Decimalization [150]. Decision [594, 98, 94, 554, 595, 37, 597]. Decisions [22]. Definitions [372]. Deming [82]. Demography [485]. Department [6, 18, 435]. Describing [529]. Descriptive [210, 355]. Design [82, 509, 559, 599, 460]. Designed [127]. Designs [493]. Development [31, 563, 275, 287]. Developments [586]. Deviates [321]. Dictionary [339]. Differences [365, 267]. Different [195]. Dilemmas [460]. Dimensions [153]. Disappearance [17]. Discrete [292]. Discriminant [76]. Discrimination [247]. Discussion [93, 348, 231]. Diseases [299]. Distribution [143, 114, 296, 209]. Distributions [557, 165]. Division [520, 41]. Doctor [16]. Does [5]. Doig [173, 499, 596]. Donald [602]. Dorothy [371]. Douglas [339]. Draper [514]. Drinking [533]. Duckworth [149]. Dunn [302].

Early [517]. Econometric [482]. Econometrics [397]. Economic [403, 207, 563, 591, 267, 561, 586, 601, 291]. Economics [576, 255, 229, 339, 320, 324, 301]. Economists [266]. Eddison [297]. Editor [236, 357, 405, 445, 500, 249]. Editorial [1, 12, 28, 40, 55, 72, 86, 102, 121, 135, 159, 414, 455]. Education [7, 21, 32, 77, 97, 112, 129, 146, 167, 186, 213, 234, 281, 235, 248, 334, 358, 408, 375, 425, 448, 467, 486, 501, 522, 538, 550, 580, 606, 124, 460]. Educational [530]. Edwards [82, 338, 292].

Effect [380]. Effectiveness [179]. Effects [379, 182, 14]. Efficiency [496]. Eighteenth [466]. Eilon [579]. Ekeblad [313]. Elandt [151]. Electric [138]. Element [588]. Elementary [464, 292, 422, 536, 600, 52, 37]. Elements [106, 352, 320, 464, 336, 497, 337, 69, 116, 154]. Emanuel [34]. Enders [81]. Engineering [347, 240, 353, 117]. Engineers [439]. England [17]. English [398, 485, 398]. English-French [398]. Environmental [496]. Epidemics [299]. Epidemiological [367]. Equipment [6]. Erratum [452, 340]. Error [491]. Errors [161]. Essentials [356]. Establishing [366]. Estimates [9]. Estimating [161, 480]. etc [108]. Ethics [495]. Evaluation [491, 380, 457, 460]. Evely [66]. Events [399]. Everyman [252]. Evidenced [380]. Examinations [426, 447, 512, 529]. Examiner [426, 447, 512, 549]. Exceptional [75]. Exchange [164, 346]. Exercises [515]. Expected [292]. Expects [16]. Experiment [476, 127, 14]. Experimental [47, 373]. Experimentation [459, 240]. Experiments [198, 373, 35, 316, 559, 599]. Explanatory [130]. Exploratory [572]. Exponential [321]. Exponentially [195]. Export [298]. Extremes [23].

F [322, 191, 482, 602, 464, 36, 115, 313, 148, 154, 189, 421]. F.R.S [144]. F.S.S [144]. Factor [3, 125, 208, 126, 123, 231, 124]. Factorial [198]. Factors [291]. Factory [113]. Family [563, 510]. Farm [31, 17]. Farmers [264]. Fatigue [316, 258]. Feller [462]. Fertility [563]. Fibre [227]. Finney [187]. Firms [374]. First [498, 115]. Fisher [144]. Fisheries [263]. Fit [209]. Fitting [321]. Flavour [532]. Food [263]. Foodstuffs [509]. Forecasting [311, 307, 452, 477, 310, 308, 478, 333, 416, 312, 393, 531, 317, 80, 318]. Forestry [558]. Foreword [262]. Formulae [372, 519].

Forrest [258]. FORTRAN [385, 294].

Forum [46]. Fothergale [345].

Foundations [175]. Fourth [190]. Fouth [190]. Fox [591, 256]. Frame [571]. Frank [103, 314]. Fred [481]. French [398, 398]. French-English [398]. Freund [464, 36]. Friday [103]. Front [2, 13, 29, 54, 71, 85, 87, 104, 122, 136, 160, 184, 196, 214, 230, 261, 271, 285, 306, 327, 344, 362, 363, 391, 413, 430, 454, 472, 490, 506, 527, 543, 553, 569, 585, 609]. Function [441]. Functions [76, 38, 296, 336]. Fundamental [320, 301]. Future [587].

G [51, 534, 579, 338, 173, 499, 596, 439, 258, 316, 497, 560, 355, 153, 422, 115, 78, 237, 519, 600, 575, 382, 217, 238, 442, 114, 219, 598, 204, 52, 336, 318]. G.P.O. [46]. Gambling [189]. Game [303]. Games [189]. Garment [557]. Geary [336]. Geese [379]. General [19, 65, 411, 410, 451, 470, 488, 503, 466, 525, 541, 551, 583, 605, 138]. Genesis [461]. Geographic [365]. Geological [254]. Geometrical [238]. Geometry [153, 432]. George [481, 219, 220]. Gerhard [397]. Glossary [398]. Godfrey [51]. Gods [189]. Godwin [296]. Goldsmith [319]. Goldstein [355]. Gotkin [355]. Gourevitch [371]. Government [555]. Graham [400]. Graphical [310]. Graybill [191]. Greater [556]. Green [441]. Greenberg [217]. Greenwald [339]. Greenwood [152]. Gregg [317]. Group [139, 274, 279, 273]. Growth [321, 374, 291, 290]. Gruenberger [481]. Guenther [383]. Guide [385, 149, 215, 113, 293, 152]. Guides [113]. Gumbel [23]. Gunnar [9]. Guttman [353, 4].

H [8, 439, 240, 317, 592, 401, 78, 319, 483, 294, 443, 25, 24, 152, 296, 113, 516, 335, 318, 514, 220, 251, 56]. Habit [368]. Habits [366]. Hacking [440]. Handbook [339, 575, 483, 218, 387, 188]. Harald [26]. Harris [593]. Hartley [8, 152]. Harvey [200]. Health [113, 394, 190]. Henrici [337]. Herman [402, 37]. Hertz [297]. Hill [339, 169, 202]. Hints [349]. Hiorns [321, 372]. Historical [485]. Hoel [422, 237, 52]. Hogg [171]. Holman [420]. Home [349]. Horse [17]. Horst [300]. Horwitz [384]. Hospital [347]. Hossell [317, 318]. Howard [597, 116]. Huitson [356, 463]. Hunter [421]. Huntsberger [320]. Hypotheses [25].

IBM [484]. Ideas [199, 400, 220]. II [563, 114, 297, 387, 462]. III [563]. Ilersic [79, 335]. Ill [113]. Ill-Health [113]. Imperial [88]. Including [404]. Incomes [3]. Incorporated [10]. Index [388]. Industrial [579, 278]. Industries [88]. Industry [350, 415, 141, 227, 142, 435, 277, 62, 225, 228, 329, 576, 66]. Inequalities [296]. Infectivity [289]. Inference [114, 438, 320, 440, 175, 383, 203]. Infinitely [81]. Influence [509]. Information [158, 224, 283, 360, 428, 504, 567, 98]. Innovation [544]. Input [31]. Input-Output [31]. Inspector [330]. Inspectorate [113]. Institute [101, 119, 133, 156, 164, 342, 411, 540, 348, 495, 229, 228]. Instruction [545, 250]. Instrument [417]. Intelligence [556]. Inter [31]. Inter-Farm [31]. Interbreed [265]. Intermediate [591]. International [520, 291]. Interplay [474]. Interplot [14]. Interpretation [31, 47, 367]. Intra [31]. Intra-Farm [31]. Introducing [603]. Introduction [403, 436, 371, 187, 473, 191, 456, 81, 99, 170, 520, 294, 123, 340, 595, 485, 384, 462, 423, 513, 131, 577, 237, 174, 315, 219, 400, 203, 562, 599]. Introductory [597, 272, 353, 210]. Intuitive [220]. Inventories [386]. Inventory [288, 579]. Investigation [161]. Investigations [245]. Investory [200]. Ira [384]. Irving [252].

Issue [123]. IV [563, 385].

J [51, 320, 420, 579, 594, 23, 187, 439, 190, 254, 374, 517, 537, 317, 464, 592, 36, 78, 398, 575, 441, 175, 352, 268, 299, 152, 296, 154, 113, 298, 335, 318, 604, 465, 578]. J. [517]. Jaffray [481]. Jagjit [444]. Japan [110]. Jean [302]. Joan [356]. John [301, 464, 218, 387, 373, 220]. Johnson [316]. Joseph [219]. Juke [417]. Jury [417].

Kahn [254]. Kalton [400]. Karl [591]. Keen [356]. Keeping [203]. Keilson [441]. Kendall [534, 173, 499, 596, 238, 114, 153]. Khintchine [48]. King [579]. Kingdom [530]. Kish [396]. Knowledge [579, 286]. Kyburg [401].

L [168, 250, 254, 577, 355, 147, 192, 319, 564, 174, 175, 294, 204, 113, 421, 335, 215, 318, 292, 465]. Lady [323]. Laha [575, 336]. Lambe [519]. Lampkin [579]. Language [369, 294]. Law [332]. Lawrence [561]. Leadbetter [516]. Leadership [264]. Learning [547]. Least [151]. Lectures [597]. Lehmann [25]. Leicester [107]. Leonard [420, 316]. Leslie [396]. Lewis [220, 399]. Life [293, 557, 144]. Limited [106, 138, 109, 164]. Lincoln [37]. Lindley [315]. Linear [191, 336, 599, 438]. Linguistics [434]. Linnik [151]. Literature [173, 499, 596]. Little [66]. Local [555]. Logic [440]. London [556, 164]. Londoner [130]. Long [333]. Longitudinal [528]. Lord [183]. Loveday [498, 68]. Ltd. [520, 576, 598]. Luck [323]. Lukacs [336, 38].

M [534, 173, 499, 596, 153, 354, 99, 170, 192, 575, 66, 382, 238, 114, 24, 171, 113, 200, 516, 336, 565]. MacDonald [439]. Machinery [598]. Machines [520]. Machol [98]. Macon [216]. Made [227]. Making [554, 192]. Makower [565]. Man [227]. Man-Made [227]. Manage [598].

Management [520, 105, 545, 200, 229]. Manager [215]. Mandel [373, 373]. Manual [130]. Manufacture [276]. Manufacturing [435]. Many [81]. Mark [192]. Market [49, 80, 67, 278, 4, 465]. Marketing [465]. Markov [594, 69]. Martin [303, 594]. Matching [333]. Materials [276]. Mathematical [190, 464, 237, 152, 116, 154, 421, 317, 354, 576, 172, 48, 253]. Mathematics [324, 515, 439, 322, 518, 604, 439]. Matrix [404, 300]. Matter [11, 27, 39, 2, 13, 29, 53, 70, 84, 54, 71, 85, 95, 120, 134, 157, 87, 104, 122, 136, 178, 194, 206, 222, 160, 184, 196, 214, 242, 260, 270, 284, 230, 261, 271, 285, 305, 326, 343, 361, 306, 327, 344, 362, 377, 390, 412, 429, 363, 391, 413, 430, 453, 471, 489, 505, 454, 472, 490, 506, 526, 542, 552, 568, 527, 543, 553, 569, 584, 608, 585, 609]. Maxwell [83]. McCarthy [336]. McCollough [250]. McCracken [385]. McGraw [339]. McGraw-Hill [339]. Measurement [264, 528, 289, 419, 496, 179]. Measurements [255, 58]. Measuring [180, 140]. Medical [243, 245, 244, 16, 169, 62, 246]. Medicine [247, 202]. Meeting [20, 19, 65, 411, 410, 466, 605]. Members [450, 468, 540, 524, 582]. Mendenhall [562, 599]. Metal [106, 109, 107]. Metallurgical [287]. Metals [258]. Meteorological [311, 452]. Meteorology [190]. Meterological [309]. Method [287, 313, 126, 251, 480, 321, 151]. Methodological [570]. Methodology [117]. Methods [371, 286, 161, 91, 99, 519, 575, 441, 308, 288, 564, 202, 442, 458, 290, 48, 90, 482, 560, 602, 600, 256, 204, 220]. Metric [33]. Migration [563]. Milk [265]. Millar [318]. Miller [254, 443]. Miner [63]. Mining [58]. Ministry [263]. Minutes [411]. Mixed [480]. Model [576, 480]. Models [207, 191, 482, 537, 479, 599, 601]. Modern [339, 464, 36, 34]. Money [437].

Monitoring [457]. Monographs [256]. Moore [338]. Moran [238]. Morgan [595]. Morrell [335, 592, 318]. Morris [255]. Morrison [602]. Mortality [368, 563, 63]. Moser [267]. Moses [37]. Mosteller [115]. Motion [350]. Motor [171]. Movements [227]. Mr. [103, 56]. Multi [161, 395]. Multi-Stage [161]. Multi-Way [395]. Multidimensional [432]. Multiple [588]. Multivariate [572, 432, 458, 602]. Murder [75]. Murray [201, 419].

N [515, 148, 352, 386, 189, 514]. Nathaniel [216]. National [57]. Natural [352]. Negative [321]. Nemchinov [324]. Nesting [379]. Neter [320, 301]. News [100, 118, 132, 155, 177, 193, 205, 221, 241, 259, 269, 304, 341, 359, 376, 389, 409, 427]. Neyman [190, 517, 190, 517]. Nigerian [171]. Nixon [398]. No [113]. Noether [497]. Noise [570]. Non [493, 176, 395, 480]. Non-Orthogonal [395, 480]. Non-Parametric [176]. Non-Standard [493]. Nonparametric [497, 218, 387]. Notes [236, 357, 405, 445, 500, 508, 249]. Notices [15]. Numbers [388]. Numerical [337, 431, 216].

O [8, 152, 402, 318]. Obituaries [407].

Obituary [103, 183]. Objective [310, 308, 333]. Observation [151]. Observations [354]. Office [42, 41]. Official [244, 398, 113]. Olive [302]. Open [46]. Opening [137, 56]. Operation [444]. Operational [149, 215, 61, 565]. Operations [168, 297]. Opler [484]. Order [217]. Organisation [57]. Organization [579, 286, 6, 88, 107, 74, 435, 507]. Organizational [74]. Organizations [339]. Orthogonal [395, 480]. Output [31]. Owen [188].

P

[338, 439, 258, 422, 319, 238, 399, 219, 52, 318].

Packaging [277]. Packed [332]. Paired [251]. Panel [507]. Paper [274, 273]. Papers [123, 517, 535]. Parametric [176]. Parcel [44]. Parr [51]. Partial [590]. Parzen [34]. Past [587]. Path [579]. Patrick [215]. Pattern [333]. Paul [237, 300]. Pay [5]. Pearce [423]. Pearson [8]. Peng [559]. Performance [392, 298]. Personal [3]. Peter [337]. Philips [139]. Physical [373]. Physics [190]. Pigeon [380]. Plan [461]. Plank [299]. Planning [563, 35]. Plant [299]. Pocket [437]. Point [329]. Pollution [364, 367]. Population [73, 563]. Populations [416, 378]. Possible [368]. Post [42, 41]. Prabhu [386]. Practical [546, 530, 124]. Practice [459, 579, 335]. Pre [332]. Pre-Packed [332]. Precision [457]. Predicting [479]. Prediction [476, 195]. Predictive [417, 590]. Predictors [474]. Preliminary [161]. Present [587]. Press [164]. Prevalence [365]. Prevaricative [232]. Preventive [202]. Primary [571]. Primer [302]. Principal [309, 127]. Principles [151, 459, 509, 78, 169, 382, 314]. Prior [94]. Probabilities [94]. Probability [190, 515, 441, 483, 315, 34, 219, 340, 462, 562, 115, 323, 401, 238, 22, 26, 115, 252]. Problem [371, 94, 162, 329]. Problem-Solving [371]. Problems [347, 328, 594, 190, 275, 570, 74, 277, 431, 61, 578, 481, 592]. Procedure [475, 381]. Procedures [78]. Proceedings [190, 190]. Processes [374, 577, 99, 352, 69, 537, 564, 201, 443, 402, 98, 595, 516]. Product [458]. Production [143, 91]. Profiles [572]. Program [531]. Programme [250, 520]. Programmed [545]. Programming [108, 518, 336, 385, 253, 484]. Progress [168, 297]. Projection [563]. Proschan [421]. Published [123]. Publishing [598].

Qualitative [83]. Quality [140, 106, 108, 573, 105, 356, 107, 110, 277, 109].

Quantitative [403, 384, 378]. Quenouille [24]. Queries [208]. Question [5]. Questions [93, 125]. Queueing [61, 48]. Queues [174, 147, 386]. Quick [212].

R [321, 388, 579, 22, 35, 49, 130, 131, 168, 254, 297, 293, 372, 404, 68, 115, 147, 78, 79, 170, 601, 575, 576, 66, 256, 319, 176, 257, 399, 443, 518, 151, 421, 516, 335, 336, 514]. Ractliffe [464]. Radhakrishna [438]. Rahman [515]. Raiffa [597]. Railway [444]. Random [219, 314, 292, 321, 374, 201, 239]. Range [333]. Rao [438]. Rapid [161, 24]. Ratcliffe [154]. Rates [368]. Rating [508]. Rats [290]. Recoveries [380]. Reed [274, 273]. Reference [78, 99, 62]. Regression [321, 589, 588, 514]. Regressor [589]. Reichmann [268, 604]. Reid [69]. Related [303, 518, 516]. Relating [332]. Relationship [165, 114]. Relationships [492]. Reliability [529, 421]. Remarks [272]. Renewal [257]. Report [410, 466, 605, 563]. Reports [426, 447, 512, 549]. Reproducibility [64]. Research [168, 297, 316, 244, 16, 91, 456, 275, 80, 556, 6, 433, 181, 530, 62, 67, 278, 4, 558, 61, 367, 403, 49, 82, 149, 215, 565]. Response [265]. Results [47]. Retrospect [366]. Review [51, 236, 302, 320, 321, 322, 323, 324, 388, 403, 420, 498, 534, 579, 594, 597, 423, 8, 22, 23, 26, 35, 49, 301, 371, 481, 338, 173, 499, 596, 187, 439, 513, 593, 130, 117, 131, 168, 190, 191, 240, 250, 254, 258, 297, 316, 374, 482, 497, 517, 535, 537, 560, 563, 577, 591, 38, 440, 515, 602, 82, 293, 355, 372, 267, 153, 317, 354, 385, 404, 422, 464, 536, 561, 592, 36, 68, 444, 401, 81, 115, 147, 78, 79, 99, 169, 170, 192, 237, 313, 396, 397, 398, 519, 520, 600, 601, 80]. Review [575, 576, 441, 484, 33, 66, 216, 256, 319, 337, 382, 483, 303, 564, 463, 148, 149, 150, 174, 175, 176, 201, 202, 217, 218, 238, 239, 257, 294, 295, 315, 352, 353, 386, 387, 399, 442, 443, 518, 268, 9, 25, 34, 69, 299, 114, 172, 24, 152, 402, 296, 219, 598, 83, 116, 154, 171, 199, 204, 255, 113, 98, 188, 189, 314, 595, 50, 151, 356, 373, 400, 421, 200, 274, 438, 52, 339, 485, 298, 516, 335, 336, 210, 215, 252, 318, 383, 384, 514, 37, 67, 419, 565, 604, 203, 292, 48, 378]. Review [220, 251, 249, 465, 253, 312, 300, 462, 603, 559, 562, 578, 599, 236, 357, 405, 445, 500, 249]. Richardson [317]. Rickmers [513]. Ringed [380]. Rivett [215]. Robert [498, 98]. Robinson [81, 239]. Roelandts [171]. Role [278]. Ronald [144]. Rosenblatt [201]. Rotherham [130]. Rourke [115]. Routine [89]. Roy [575]. Russell [215].

S [324, 579, 423, 8, 254, 535, 355, 404, 99, 353, 172, 298, 565, 203, 253]. S. [535]. Sales [477, 393, 531]. Salzman [561]. Sampford [170]. Sample [82]. Samples [161]. Sampling [76, 536, 170, 571, 45, 44, 396, 199]. Samuel [172]. Sarhan [217]. Savage [175, 176]. Sc.D. [144]. Scale [4]. Scales [508]. Scaling [64, 67]. Scheduling [579]. Schemes [143]. Schenck [240]. Schlaifer [22]. Schumaker [220]. Science [187, 117, 33]. Sciences [404, 302, 371, 254, 78, 352]. Scientific [199, 559]. Scientist [603]. Scientists [439, 400, 300]. Scope [88]. Scott [267]. Search [58]. Searle [404]. Seasonal [475, 479]. Second [68]. Section [281]. Security [113]. Selection [517]. Self [250]. Self-Instruction [250]. Semi [392]. Semi-Continuous [392]. Seminar [593]. Senior [394]. Sensory [507, 509]. Sequential [442]. Series [311, 452, 476, 399, 402, 479, 593]. Service [18]. Seventeenth [410]. Sewell [465]. Shchigolev [354]. Shirley [229, 228]. Shooting [380]. Short [393, 531, 318]. Short-Term [393, 531, 318]. Should [33]. Shubik [303]. Sick [42]. Signals [474]. Significance [370]. Significant [58].

Simple [369, 94]. Simulation [415, 295]. Since [586]. Singh [444]. Sir [144]. Skewed [143]. Small [18]. Smallwood [294]. Smith [147, 514]. Smoking [368, 364, 366]. Smokler [401]. Snedecor [560]. Social [537, 267, 303, 113, 400, 300, 603]. Society [80, 67]. Solution [481]. Solving [371, 578]. Some [381, 275, 125, 208, 586, 508, 290, 181, 274, 74, 182, 277, 530, 479, 431, 61, 393, 263]. Sometimes [30]. Source [368]. Sources [113]. Special [78, 99, 62]. Specifications [458]. Spectral [593, 475]. Spell [604]. Spitzer [314]. Square [209]. Squares [151]. Staff [484, 394]. Stage [161]. Standard [161, 493, 338]. Standards [331]. Station [6]. Stationary [516]. Statistic [209]. Statistical [347, 320, 207, 338, 173, 499, 596, 187, 364, 191, 316, 286, 517, 381, 440, 602, 587, 91, 57, 106, 267, 275, 287, 115, 313, 600, 308, 256, 288, 175, 399, 25, 6, 24, 18, 138, 88, 126, 204, 188, 595, 227, 373, 400, 142, 274, 438, 380, 74, 276, 435, 383, 203, 431, 509, 312, 90, 371, 117, 250, 254, 560, 293, 444, 519, 202, 9, 255, 50, 200, 220]. Statistically [310]. Statistician [16, 369, 329, 41]. Statisticians [20, 8, 30, 10, 101, 119, 133, 156, 342, 411]. Statistics [302, 320, 331, 328, 243, 423, 22, 26, 301, 92, 266, 245, 494, 547, 439, 190, 244, 515, 307, 456, 141, 546, 404, 586, 78, 398, 226, 315, 387, 573, 114, 545, 598, 139, 204, 113, 340, 89, 555, 42, 252, 62, 225, 246, 548, 228, 603, 578, 263, 464, 335, 220, 498, 534, 190, 497, 591, 355, 372, 422, 464, 36, 68, 169, 237, 520, 575, 382, 483, 176, 217, 218, 353, 442, 268, 172, 152, 116, 154, 113, 52, 292, 603, 562, 420, 23, 371, 513, 79, 192, 335, 210, 603]. Statistics. [403]. Status [449, 469, 487, 502, 523, 539, 566, 581, 607]. Steel [78]. Steindl [374]. Steven [321]. Stochastic [577, 99, 564, 352, 443, 402, 516, 459, 537]. Stock [346, 227]. Stocks [592]. Storage [592]. Strategy [511]. Stratification [571].

Stratified [161]. Structure [273]. Stuart [534, 114, 199]. Student [408, 349, 348]. Students [371, 372]. Studies [528, 401]. Study [364, 267, 365, 433, 373]. Subject [479]. Subjective [401, 494]. Success [379]. Suggested [395]. Suits [403]. Sum [319]. Summary [93, 563, 231]. Sums [590]. Superior [350]. Supervisors [140]. Supply [265]. Survey [491, 437, 290, 570, 418, 533, 396]. Surveys [572]. Symposium [190]. Syski [131]. System [392, 33, 484]. System/360 [484]. Systems [195, 131, 239, 200].

T [297, 535, 317, 352, 69]. Table [370, 395]. Tables [8, 483, 152, 188]. Takacs [564, 174]. Taro [536]. Taxonomic [433]. Taxonomy [431]. Teaching [266, 547, 141, 573, 545, 139, 544, 548]. Technical [51, 548]. Technique [316, 310, 395]. Techniques [49, 572, 18, 393, 319]. Technological [291, 419]. Technologists [439]. Technology [229]. Telephone [131, 45]. Ten [460]. Term [393, 531, 318]. Terms [398, 339]. Test [209]. Testing [43, 163, 370, 458, 507, 25]. Tests [290, 509, 532]. Texts [256]. Their [69, 267]. Theoretical [392]. Theories [240]. Theory [323, 388, 534, 345, 117, 131, 190, 303, 564, 174, 443, 34, 69, 114, 151, 421, 335, 419, 48, 462, 190, 536, 170, 441, 257, 37, 316]. Thirkettle [204]. Thomas [115, 115]. Thornley [579]. Tickets [45]. Time [593, 311, 452, 476, 402, 479]. Tintner [397]. Tocher [295]. Top [212]. Topic [510]. Torrie [78]. Total [491]. Towns [267]. Toxicological [290]. Tracking [474]. Traffic [171]. Training [350, 226, 142, 109]. Transformed [9]. Treatment [316, 94]. Trend [317]. Trends [563]. Trials [557]. Tutor [520]. Twentieth [605]. Two [572, 312]. Type [321].

U [386]. Uncertainty [597].

Undergraduates [515]. Unit [556]. United [530]. Units [571]. Upon [379]. Urbanization [563]. Usage [484]. Use [324, 279, 76, 244, 311, 452, 601, 108, 232, 478, 268, 64, 555, 4, 246, 367]. Used [49, 18]. Uses [207, 263]. Using [94].

V [320, 321, 324, 579, 371, 317, 315, 518, 171, 151]. Vajda [253]. Value [124, 590]. Values [292]. Variable [143]. Variables [589, 292, 9, 219]. Variance [480, 463]. Variates [81]. Variation [316, 433, 479]. Various [368]. Vehicle [171]. Vernon [220]. View [329]. Viewpoint [330, 315]. Virus [289]. Visit [164]. Vol [168, 297, 387, 462, 603]. Volume [158, 224, 283, 360, 428, 504, 567, 8, 190, 191, 563, 114, 26].

W [51, 320, 321, 388, 579, 49, 535, 560, 82, 293, 372, 398, 256, 399, 268, 116, 171, 595, 318, 604, 462, 562, 599]. Wadsworth [219]. Wagner [200]. Wales [17]. Walk [314]. Walsh [218, 387]. Walter [147]. Warren [323]. Wasserman [320, 301]. Wavelets [239]. Way [395]. Ways [212]. Wear [557]. Wear-Trials [557]. Weather [379, 307, 308]. Weaver [323]. Weighing [332]. Weight [328, 330, 329]. Weighted [195]. Weinberg [220]. Wells [298]. Wetherill [600, 442]. Wheldon [204]. Wilks [535, 353, 172]. William [301, 483, 383]. Williams [36, 113]. Williamson [565]. Within [350, 141, 109, 263]. Wold [402]. Wolf [267]. Wood [380]. Wood-Pigeon [380]. Woodward [319]. Work [57, 88, 144, 378, 113]. Working [60]. World [563, 563]. Wrigley [485]. Writer [51]. Wynn [56].

Yamane [536]. Year [410, 466, 605]. Yeomans [603]. Yu [151].

References

Anonymous:1960:Ea

[1] Anonymous. Editorial. The Incorporated Statistician, 10(1):1, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987481.

Anonymous:1960:FMa

[2] Anonymous. Front matter. The Incorporated Statistician, 10(1):2–36, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987480.

Burton:1960:RCP

[3] Harry Burton. Recent changes in personal and factor incomes. The Incorporated Statistician, 10(1):3–14, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987482.

Smith:1960:UGS

[4] John S. Smith. The use of the Guttman Scale in market research. The Incorporated Statistician, 10(1):15–28, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987483.

Kent:1960:AAQ

[5] A. J. Kent and C. Nuttall. An attempt to answer the question — does advertising pay? The Incorporated Statistician, 10(1):29–35, February 1960. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2987484.

Jeffers:1960:OES

[6] J. N. R. Jeffers. Organization and equipment of a statistical department in a biological research station. The Incorporated Statistician, 10(1):37–48, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987485.

Anonymous:1960:Eb

[7] Anonymous. Education. The Incorporated Statistician, 10(1):49–52, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987486.

B:1960:BRBa

[8] W. R. B. Book review: BIOMETRIKA Tables for Statisticians, Volume I, by E. S. Pearson; H. O. Hartley. The Incorporated Statistician, 10(1):52–53, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987487.

J:1960:BRBa

[9] N. L. J. Book review: Statistical Estimates and Transformed Beta Variables, by Gunnar Blom. The Incorporated Statistician, 10(1):53–55, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987488.

Anonymous:1960:AIS

[10] Anonymous. The Association of Incorporated Statisticians. The Incorporated Statistician, 10(1):56, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987489.

Anonymous:1960:BMa

[11] Anonymous. Back matter. The Incorporated Statistician, 10(1):??, February 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987490.

Anonymous:1960:Ec

[12] Anonymous. Editorial. The Incorporated Statistician, 10(2):57, July 1960. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2987407.

Anonymous:1960:FMb

[13] Anonymous. Front matter. The Incorporated Statistician, 10(2):58, July 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987406.

Roberts:1960:AEI

[14] P. Roberts. The analysis of an experiment on interplot effects. The Incorporated Statistician, 10(2):59–65, July 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987408.

Anonymous:1960:N

[15] Anonymous. Notices. The Incorporated Statistician, 10(2):66, July 1960. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2987409.

Chapman:1960:WDE

[16] P. J. Chapman. What the doctor expects from the statistician in medical research. The Incorporated Statistician, 10(2):67– 78, July 1960. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987410.

Britton:1960:DFH

[17] D. K. Britton. The disappearance of the farm horse in England and Wales. The Incorporated Statistician, 10(2):79– 88, July 1960. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987411.

Keen:1960:STU

[18] J. Keen. Statistical techniques used in a small statistical service department. The Incorporated Statistician, 10(2):89– 91, July 1960. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2987412.

Anonymous:1960:AGM

[19] Anonymous. Annual General Meeting. The Incorporated Statistician, 10(2):92– 93, July 1960. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2987413.

Anonymous:1960:MS

[20] Anonymous. Meeting of statisticians. The Incorporated Statistician, 10(2):93, July 1960. CODEN ???? ISSN 1466-

                                        9404. URL http://www.jstor.org/
                                        stable/2987414.

Anonymous:1960:Ed

[21] Anonymous. Education. The Incorporated Statistician, 10(2):94, July 1960. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2987415.

B:1960:BRBb

[22] W. R. B. Book review: Probability and Statistics for Business Decisions, by R. Schlaifer. The Incorporated Statistician, 10(2):95–96, July 1960. CODEN ???? ISSN 1466-9404. URL http://

                                        www.jstor.org/stable/2987416.

B:1960:BRBc

[23] W. R. B. Book review: Statistics of Extremes, by E. J. Gumbel. The Incorporated Statistician, 10(2):96–97, July 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987417.

K:1960:BRB

[24] J. K. Book review: Rapid Statistical Calculations, by M. H. Quenouille. The Incorporated Statistician, 10(2):97– 98, July 1960. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987418.

J:1960:BRBb

[25] N. L. J. Book review: Testing Statistical Hypotheses, by E. H. Lehmann. The Incorporated Statistician, 10(2):98– 99, July 1960. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987419.

B:1960:BRBd

[26] W. R. B. Book review: Probability and Statistics: The Harald Cramer Volume. The Incorporated Statistician, 10(2):99– 100, July 1960. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987420.

Anonymous:1960:BMb

[27] Anonymous. Back matter. The Incorporated Statistician, 10(2):??, July 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987421.

Anonymous:1960:Ee

[28] Anonymous. Editorial. The Incorporated Statistician, 10(3):101, October 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987469.

Anonymous:1960:FMc

[29] Anonymous. Front matter. The Incorporated Statistician, 10(3):102, October 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987468.

Beveridge:1960:SSC

[30] Lord Beveridge. Statisticians sometimes count. The Incorporated Statistician, 10(3):103–119, October 1960. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987470.

Bessell:1960:DIF

[31] J. E. Bessell. A development of an inter-farm input-output analysis for intra-farm interpretation. The Incorporated Statistician, 10(3):120–134, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987471.

Anonymous:1960:Ef

[32] Anonymous. Education. The Incorporated Statistician, 10(3):135–139, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987472.

H:1960:BRB

[33] N. K. H. Book review: Decimal Coinage and the Metric System. Should Britain Change?, by The British Association for the Advancement of Science; The Association of British Chambers of Commerce. The Incorporated Statistician, 10(3):140–141, October 1960. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987473.

J:1960:BRBc

[34] N. L. J. Book review: Modern Probability Theory and Its Applications, by Emanuel Parzen. The Incorporated Statistician, 10(3):141–143, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987474.

B:1960:BRBe

[35] W. R. B. Book review: Planning of Experiments, by D. R. Cox. The Incorporated Statistician, 10(3):143–144, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987475.

F:1960:BRB

[36] R. A. F. Book review: Modern Business Statistics, by J. E. Freund; F. J. Williams. The Incorporated Statistician, 10(3):145, October 1960. CODEN ???? ISSN 1466-9404. URL http://www.

jstor.org/stable/2987476.

S:1960:BRB

[37] C. A. B. S. Book review: Elementary Decision Theory, by Herman Chernoff; Lincoln E. Moses. The Incorporated Statistician, 10(3):146–147, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987477.

C:1960:BRB

[38] D. R. C. Book review: Characteristic Functions, by E. Lukacs. The Incorporated Statistician, 10(3):147–148, October 1960. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987478.

Anonymous:1960:BMc

[39] Anonymous. Back matter. The Incorporated Statistician, 10(3):??, October 1960. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987479.

Anonymous:1961:Ea

[40] Anonymous. Editorial. The Incorporated Statistician, 11(1):1, January 1961. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2987440.

Wood:1961:POC

[41] S. Wood. The Post Office and The Chief Statistician’s Division. The Incorporated Statistician, 11(1):2–8, January 1961. CODEN ???? ISSN 1466-9404.

                                        URL http://www.jstor.org/stable/
                                        2987441.

Perkins:1961:SAS

[42] N. A. Perkins. Sick absence statistics in the post Office. The Incorporated Statistician, 11(1):8–13, January 1961. CO-DEN ???? ISSN 1466-9404. URL http:

                                        //www.jstor.org/stable/2987442.

Balcombe:1961:AT

[43] J. Balcombe. Acceptance testing. The Incorporated Statistician, 11(1):13–17, January 1961. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2987443.

Neifield:1961:PS

[44] M. L. M. Neifield. Parcel sampling. The Incorporated Statistician, 11(1):18– 23, January 1961. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2987444.

Lane:1961:STT

[45] P. J. Lane. Sampling telephone tickets. The Incorporated Statistician, 11(1):23– 29, January 1961. CODEN ???? ISSN

                                        1466-9404. URL http://www.jstor.
                                        org/stable/2987445.

Anonymous:1961:GPC

[46] Anonymous. A.I.S./G.P.O. Conference open forum. The Incorporated Statistician, 11(1):29–32, January 1961. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2987446.

Freeman:1961:AIE

[47] G. H. Freeman. The analysis and interpretation of experimental results. The Incorporated Statistician, 11(1):33–56, January 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987447.

Syski:1961:BRB

[48] R. Syski. Book review: Mathematical Methods in the Theory of Queueing, by A. Y. Khintchine. The Incorporated Statistician, 11(1):57–61, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987448.

B:1961:BRB

[49] W. R. B. Book review: The Adequacy of Techniques Used in Market Research, by W. Belson; C. R. Bell. The Incorporated Statistician, 11(1):61–62, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987449.

M:1961:BRB

[50] M. J. M. Book review: Statistical Analysis, by E. C. Bryant. The Incorporated Statistician, 11(1):62–63, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987450.

Anonymous:1961:BRB

[51] Anonymous. Book review: The Technical Writer, by J. W. Godfrey; G. Parr. The Incorporated Statistician, 11(1):63– 64, January 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987451.

Morrell:1961:BRB

[52] A. J. H. Morrell. Book review: Elementary Statistics, by P. G. Hoel. The Incorporated Statistician, 11(1):64, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987452.

Anonymous:1961:BMa

[53] Anonymous. Back matter. The Incorporated Statistician, 11(1):??, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987453.

Anonymous:1961:FMa

[54] Anonymous. Front matter. The Incorporated Statistician, 11(1):??, January 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987439.

Anonymous:1961:Eb

[55] Anonymous. Editorial. The Incorporated Statistician, 11(2):65, September 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987423.

Wynn:1961:OAM

[56] A. H. A. Wynn. Opening address by Mr. A. H. A. Wynn. The Incorporated Statistician, 11(2):66–68, September 1961. CODEN ???? ISSN 1466-

9404. URL http://www.jstor.org/
stable/2987424.

Cowe:1961:ONC

[57] R. G. C. Cowe. The organisation of the national coal board and its statistical work. The Incorporated Statistician, 11(2):68–70, September 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987425.

Schumacher:1961:SSM

[58] E. F. Schumacher. The search for significant measurements in coal mining. The Incorporated Statistician, 11(2):70– 75, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987426.

Wakefield:1961:CC

[59] W. B. Wakefield. Colliery concentration. The Incorporated Statistician, 11(2):75– 83, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987427.

Nattrass:1961:CCW

[60] J. G. Nattrass. Concentration of colliery working. The Incorporated Statistician, 11(2):84–90, September 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987428.

Tomlinson:1961:SQP

[61] R. C. Tomlinson. Some queueing problems in operational research. The Incorporated Statistician, 11(2):90–97, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987429.

Rogan:1961:MRC

[62] J. M. Rogan. Medical research in the coal industry with special reference to statistics. The Incorporated Statistician, 11(2):98–103, September 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987430.

Liddell:1961:MM

[63] F. D. K. Liddell. Miners’ mortality. The Incorporated Statistician, 11(2):103–109, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987431.

Jobling:1961:UCR

[64] D. Jobling and E. J. Snell. The use of the coefficient of reproducibility in attitude scaling. The Incorporated Statistician, 11(2):110–118, September 1961. CO-DEN ???? ISSN 1466-9404. URL http:

//www.jstor.org/stable/2987432.

Anonymous:1961:AGM

[65] Anonymous. Annual General Meeting. The Incorporated Statistician, 11 (2):119–121, September 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987433.

H:1961:BRB

[66] J. H. Book review: Concentration in British Industry, by R. Evely; I. M. D. Little. The Incorporated Statistician, 11 (2):122–123, September 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987434.

S:1961:BRB

[67] J. S. S. Book review: Attitude Scaling, by Market Research Society. The Incorporated Statistician, 11(2):123–125, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987435.

F:1961:BRB

[68] R. A. F. Book review: A Second Course in Statistics, by R. Loveday. The Incorporated Statistician, 11(2):125–126, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987436.

J:1961:BRB

[69] G. J. Book review: Elements of the Theory of Markov Processes and Their Applications, by A. T. Bharucha-Reid. The Incorporated Statistician, 11(2):126–128, September 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987437.

Anonymous:1961:BMb

[70] Anonymous. Back matter. The Incorporated Statistician, 11(2):??, September 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987438.

Anonymous:1961:FMb

[71] Anonymous. Front matter. The Incorporated Statistician, 11(2):??, September 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987422.

Anonymous:1961:Ec

[72] Anonymous. Editorial. The Incorporated Statistician, 11(3):129, December 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987455.

Benjamin:1961:CP

[73] B. Benjamin. The 1961 census of population. The Incorporated Statistician, 11 (3):130–143, December 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987456.

Newman:1961:SOP

[74] Dennis Newman. Some organizational problems arising in a statistical organization in changing to a computer. The Incorporated Statistician, 11(3):144–149, December 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987457.

Little:1961:MEC

[75] A. N. Little. Murder: The exceptional crime. The Incorporated Statistician, 11 (3):150–155, December 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987458.

Calder:1961:UDF

[76] A. B. Calder. The use of discriminant functions in biological sampling. The Incorporated Statistician, 11(3):156–163, December 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987459.

Anonymous:1961:Ed

[77] Anonymous. Education. The Incorporated Statistician, 11(3):164–169, December 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987460.

Fox:1961:BRBa

[78] R. A. Fox. Book review: Principles and Procedures of Statistics with Special Reference to the Biological Sciences, by R. G. D. Steel; J. H. Torrie. The Incorporated Statistician, 11(3):170–171, December 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987461.

Fox:1961:BRBb

[79] R. A. Fox. Book review: Statistics, by A. R. Ilersic. The Incorporated Statistician, 11(3):171–172, December 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987462.

Friday:1961:BRB

[80] F. A. Friday. Book review: Business Forecasting, by Market Research Society. The Incorporated Statistician, 11 (3):172–173, December 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987463.

Foster:1961:BRB

[81] F. G. Foster. Book review: An Introduction to Infinitely Many Variates, by Enders A. Robinson. The Incorporated Statistician, 11(3):173, December 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987464.

Corlett:1961:BRB

[82] T. Corlett. Book review: Sample Design in Business Research, by W. Edwards Deming. The Incorporated Statistician, 11(3):173–175, December 1961. CODEN ???? ISSN 1466-9404. URL http://

www.jstor.org/stable/2987465.

Lewis:1961:BRB

[83] T. W. Lewis. Book review: Analysing Qualitative Data, by A. E. Maxwell. The Incorporated Statistician, 11(3):175–176, December 1961. CODEN ???? ISSN

1466-9404. URL http://www.jstor.
org/stable/2987466.

Anonymous:1961:BMc

[84] Anonymous. Back matter. The Incorporated Statistician, 11(3):??, December 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987467.

Anonymous:1961:FMc

[85] Anonymous. Front matter. The Incorporated Statistician, 11(3):??, December 1961. CODEN ???? ISSN 1466-9404.

URL http://www.jstor.org/stable/
2987454.

Anonymous:1962:Ea

[86] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):1, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987222.

Anonymous:1962:FMa

[87] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):2–3, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987221.

Kenney:1962:SWI

[88] H. Kenney. Statistical work in Imperial chemical industries: Its scope and organization. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):4–9, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987223.

Mann:1962:RAC

[89] D. F. C. Mann. Routine administrative and commercial statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):10–15, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987224.

Welford:1962:CAS

[90] B. P. Welford. Commercial applications of statistical methods. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):16–22, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987225.

Coutie:1962:SMR

[91] G. A. Coutie. Statistical methods in research and production. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):23–31, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987226.

Baines:1962:SC

[92] A. Baines. Statistics and computers. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1): 32–38, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987227.

Anonymous:1962:DSQ

[93] Anonymous. Discussion: Summary of questions and discussion. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):39–42, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987228.

Lindley:1962:TSD

[94] D. V. Lindley. A treatment of a simple decision problem using prior probabilities. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (1):43–53, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987229.

Anonymous:1962:BMa

[95] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):54–60, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987236.

Smith:1962:Ca

[96] John S. Smith and D. V. Lindley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):55–56, January 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987230.

Anonymous:1962:Eb

[97] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):57, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987231.

Lindley:1962:BRB

[98] D. V. Lindley. Book review: Information and Decision Processes, by Robert E. Machol. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):58, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987232.

Fox:1962:BRB

[99] R. A. Fox. Book review: An Introduction to Stochastic Processes with Special Reference to Methods and Applications, by M. S. Bartlett. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):59, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987233.

Anonymous:1962:NAa

[100] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 12(1): 61–66, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987234.

Anonymous:1962:ISa

[101] Anonymous. The Institute of Statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 12(1):67, January 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987235.

Anonymous:1962:Ec

[102] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):69–70, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987054.

Anonymous:1962:OMF

[103] Anonymous. Obituary: Mr. Frank A.

Friday. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):70, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987055.

Anonymous:1962:FMb

[104] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):71–132, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987053.

Kent:1962:MAQ

[105] A. N. Kent. The management approach to quality control. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):72–75, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987056.

Daisley:1962:SEQ

[106] P. A. Daisley. The statistical elements of quality control in the metal Box company limited. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):76–89, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987057.

Moy:1962:OAQ

[107] A. G. Moy. The organization and application of quality control at Metal Box, Leicester. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):90–101, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987058.

Gearing:1962:UAP

[108] H. W. Gearing. Use of autocodes for programming calculations for quality control, etc. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):102–110, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987059.

Slight:1962:TQC

[109] P. F. Slight. Training in quality control: Within the metal Box company limited. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):111–117, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987060.

Nixon:1962:QAJ

[110] Frank Nixon. Quality achievements in Japan. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (2):119–124, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987061.

Smith:1962:Cb

[111] C. F. Smith, Joan Keen, and D. V. Lindley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):125–126, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987062.

Anonymous:1962:Ed

[112] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):129–131, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987063.

Liddell:1962:BRB

[113] F. D. K. Liddell. Book review: Guides to Official Sources: No. 5. Social Security Statistics; Guide to Statistics Collected by H. M. Factory Inspectorate; Accidents and Ill-Health at Work, by J. L. Williams. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):133–138, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987064.

Johnson:1962:BRB

[114] N. L. Johnson. Book review: The Advanced Theory of Statistics: Volume I (1958): Distribution Theory; The Advanced Theory of Statistics. Volume II (1961): Inference and Relationship, by M. G. Kendall; A. Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):138–142, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987065.

Foster:1962:BRBa

[115] F. G. Foster. Book review: Probability: A First Course, by F. Mosteller; R. E. K. Rourke; G. B. Thomas; Probability With Statistical Applications, by F. Mosteller; R. E. K. Rourke; G. B. Thomas. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):142–143, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987066.

Lewis:1962:BRBa

[116] T. W. Lewis. Book review: Elements of Mathematical Statistics, by Howard W. Alexander. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):143–144, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987067.

Buckland:1962:BRBa

[117] W. R. Buckland. Book review: Statistical Theory and Methodology in Science and Engineering, by K. A. Brownlee. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (2):144–145, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987068.

Anonymous:1962:NAb

[118] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 12(2): 146–153, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987069.

Anonymous:1962:ISb

[119] Anonymous. The Institute of Statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (2):154–164, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987070.

Anonymous:1962:BMb

[120] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(2):??, ???? 1962.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987071.

Anonymous:1962:Ee

[121] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):165, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986911.

Anonymous:1962:FMc

[122] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):166–256, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986910.

Lindley:1962:FAI

[123] D. V. Lindley. Factor analysis: An introduction to the papers published in this issue. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (3):169–171, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986912.

Warburton:1962:PVF

[124] F. W. Warburton. The practical value of factor analysis in education. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):172–188, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986913.

Ehrenberg:1962:SQA

[125] A. S. C. Ehrenberg. Some questions about factor analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):191–208, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986914.

Lawley:1962:FAS

[126] D. N. Lawley and A. E. Maxwell. Factor analysis as a statistical method. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3): 209–229, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986915.

Jeffers:1962:PCA

[127] J. N. R. Jeffers. Principal component analysis of designed experiment. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3): 230–242, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986916.

Neumark:1962:C

[128] Z. Neumark, R. V. Brown, Mrs. Knight, and Joan Keen. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):243–245, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986917.

Anonymous:1962:Ef

[129] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):247–249, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986918.

Broadbent:1962:BRB

[130] S. R. Broadbent. Book review: Explanatory Manual for the Londoner, by D. R. Rotherham. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):251–254, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986919.

Buckland:1962:BRBb

[131] W. R. Buckland. Book review: Introduction to Congestion Theory in Telephone Systems, by R. Syski. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):254–255, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986920.

Anonymous:1962:NAc

[132] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3): 257–264, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986921.

Anonymous:1962:ISc

[133] Anonymous. The Institute of Statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (3):265–266, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986922.

Anonymous:1962:BMc

[134] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(3):267–268, ????

1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986923.

Anonymous:1962:Eg

[135] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):269, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986943.

Anonymous:1962:FMd

[136] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):271–350, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986942.

Thomas:1962:OA

[137] L. A. Thomas and O. W. Humphries.

Opening address. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):273–275, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986944.

Keen:1962:SCG

[138] Joan Keen and A. N. James. Statistical classes at the general electric company limited. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):276–280, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986945.

Lesser:1962:TSP

[139] R. H. S. Lesser. The teaching of statistics in the Philips group. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):281–285, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986946.

Daisley:1962:MQC

[140] P. A. Daisley. “Measuring quality” — courses for supervisors. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):286–288, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986947.

Cunliffe:1962:TSW

[141] S. V. Cunliffe. The teaching of statistics within industry … a comment. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4): 289–291, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986948.

Moore:1962:STI

[142] P. G. Moore. Statistical training on an industry basis. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):292–295, ???? 1962. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986949.

Burrows:1962:CSP

[143] Peter M. Burrows. X̄ control schemes for a production variable with skewed distribution. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (4):296–312, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986950.

Owen:1962:ALW

[144] A. R. G. Owen. An appreciation of the life and work of Sir Ronald Aylmer Fisher: F.R.S., F.S.S., Sc.D. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4): 313–319, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986951.

Bell:1962:C

[145] Roland Bell, Joan Keen, and G. A.

Barnard. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):321–325, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986952.

Anonymous:1962:Eh

[146] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):327–330, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986953.

Foster:1962:BRBb

[147] F. G. Foster. Book review: Queues, by D. R. Cox; Walter L. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):332, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986954.

Huitson:1962:BRBa

[148] A. Huitson. Book review: Combinatorial Chance, by F. N. David; D. E. Barton. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):333, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986955.

Huitson:1962:BRBb

[149] A. Huitson. Book review: A Guide to Operational Research, by E. Duckworth. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (4):333–334, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986956.

Huitson:1962:BRBc

[150] A. Huitson. Book review: The Case Against Decimalization, by A. C. Aitken. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):334–335, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986957.

MacDonald:1962:BRB

[151] P. MacDonald. Book review: Method of Least Squares and Principles of the Theory of Observation, by Yu. V. Linnik; R. C. Elandt. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):335–336, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986958.

Keen:1962:BRB

[152] Joan Keen. Book review: Guide to Tables in Mathematical Statistics, by J. Arthur Greenwood; H. O. Hartley. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4): 336–337, ???? 1962. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986959.

Draper:1962:BRB

[153] J. Draper. Book review: A Course in the Geometry of n-Dimensions, by M. G. Kendall. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):337–338, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986960.

Lewis:1962:BRBb

[154] T. W. Lewis. Book review: Elements of Mathematical Statistics, by J. F. Ratcliffe. Journal of the Royal Statistical Society. Series D (The Statistician), 12 (4):338–339, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986961.

Anonymous:1962:NAd

[155] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 12(4): 341–349, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986962.

Anonymous:1962:ISd

[156] Anonymous. The Institute of Statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):351, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986963.

Anonymous:1962:BMd

[157] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):352, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986964.

Anonymous:1962:VI

[158] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 12(4):353–356, ???? 1962. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986941.

Anonymous:1963:Ea

[159] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):1, ???? 1963. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986658.

Anonymous:1963:FMa

[160] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):2–78, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986657.

Corlett:1963:RME

[161] T. Corlett. Rapid methods of estimating standard errors of stratified multistage samples: A preliminary investigation. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):5–16, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986659.

Smith:1963:PA

[162] J. S. Smith. Problem analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):17–32, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986660.

Brook:1963:TC

[163] Caspar Brook. Testing by consumers.

Journal of the Royal Statistical Society. Series D (The Statistician), 13(1): 33–43, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986661.

Anonymous:1963:VIL

[164] Anonymous. Visit of The Institute to the London Press Exchange Limited. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (1):47–53, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986662.

Jowett:1963:RBB

[165] G. H. Jowett. The relationship between the binomial and F distributions. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (1):55–57, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986663.

Liddell:1963:C

[166] F. D. K. Liddell, D. V. Lindley, and A. A. Dewey. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):59–62, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986664.

Anonymous:1963:Eb

[167] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):65–77, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986665.

Buckland:1963:BRBa

[168] W. R. Buckland. Book review: Progress in Operations Research: Vol. 1, by R. L. Ackoff. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):79–80, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986666.

Fox:1963:BRBa

[169] R. H. Fox. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):80, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986667.

Fox:1963:BRBb

[170] R. A. Fox. Book review: An Introduction to Sampling Theory, by M. R. Sampford. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):81–82, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986668.

Lewis:1963:BRBa

[171] T. W. Lewis. Book review: Nigerian Motor Vehicle Traffic, by V. W. Hogg; C. M. Roelandts. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):82–83, ???? 1963. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986669.

Johnson:1963:BRB

[172] N. L. Johnson. Book review: Mathematical Statistics, by Samuel S. Wilks. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1): 83–84, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986670.

Barton:1963:BRB

[173] D. E. Barton. Book review: Bibliography of Statistical Literature, 1950–1958, by M. G. Kendall; Alison G. Doig. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):84, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986671.

Huitson:1963:BRBa

[174] A. Huitson. Book review: Introduction to the Theory of Queues, by L. Takacs. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):85, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986672.

Huitson:1963:BRBb

[175] A. Huitson. Book review: The Foundations of Statistical Inference, by L. J.

Savage. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):85–86, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986673.

Huitson:1963:BRBc

[176] A. Huitson. Book review: Bibliography on Non-Parametric Statistics, by I. R. Savage. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):86, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986674.

Anonymous:1963:NAa

[177] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1): 87–96, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986675.

Anonymous:1963:BMa

[178] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(1):97–100, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986676.

Treasure:1963:MAE

[179] J. A. P. Treasure. The measurement of advertising effectiveness. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):101–102, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986966.

Joyce:1963:AMA

[180] Timothy Joyce. An approach to measuring the achievement of advertising. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2): 103–116, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986967.

Monk:1963:SAA

[181] D. M. Monk. Some aspects of advertising research. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):117–126, ???? 1963. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986968.

Pridmore:1963:SEA

[182] W. A. Pridmore and E. J. Snell. Some effects of advertising. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):127–136, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986969.

George:1963:OLB

[183] C. O. George. Obituary: Lord Beveridge. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (2):137–143, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986970.

Anonymous:1963:FMb

[184] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):144–146, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986965.

Umoh:1963:C

[185] D. E. Umoh and D. V. Lindley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):147–148, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986971.

Anonymous:1963:Ec

[186] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):149–155, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986972.

Bessell:1963:BRB

[187] J. E. Bessell. Book review: An Introduction to Statistical Science in Agriculture, by D. J. Finney. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):156–157, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986973.

Lindley:1963:BRBa

[188] D. V. Lindley. Book review: Handbook of Statistical Tables, by D. B. Owen. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2): 157–158, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986974.

Lindley:1963:BRBb

[189] D. V. Lindley. Book review: Games, Gods and Gambling, by F. N. David.

Journal of the Royal Statistical Society. Series D (The Statistician), 13(2): 158–159, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986975.

Buckland:1963:BRBb

[190] W. R. Buckland. Book review: Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Volume 1: Theory of Statistics, by J. Neyman; Proceedings of the Fouth Berkeley Symposium on Mathematical Statistics and Probability. Volume 2: Probability Theory, by J. Neyman; Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability. Volume 3: Astronomy, Meteorology and Physics, by J. Neyman; Proceedings of the Fourth Berkeley Symposium on Mathematical Statistics and Probability: Volume 4: Biology and Problems of Health, by J. Neyman. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (2):159–160, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986976.

Buckland:1963:BRBc

[191] W. R. Buckland. Book review: An Introduction to Linear Statistical Models. Volume 1, by F. A. Graybill. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2): 160–161, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986977.

Fox:1963:BRBc

[192] R. A. Fox. Book review: Statistics in the Making, by M. L. Mark. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):161–162, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986978.

Anonymous:1963:NAb

[193] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2): 163–169, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986979.

Anonymous:1963:BMb

[194] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(2):170–172, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986980.

Ward:1963:CDS

[195] Denis H. Ward. Comparison of different systems of exponentially weighted prediction. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):173–185 + 187–191, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986810.

Anonymous:1963:FMc

[196] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):192–235, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986809.

Hill:1963:C

[197] T. Hill, D. V. Lindley, G. A. R. W.

Taylor, D. V. Lindley, W. A. Clough, D. J. Bartholomew, Sigma, P. Macdonald, and J. E. Bessell. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3): 193–202, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986811.

James:1963:FE

[198] A. N. James. Factorial experiments.

Journal of the Royal Statistical Society. Series D (The Statistician), 13(3): 203–209, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986812.

Lewis:1963:BRBb

[199] T. W. Lewis. Book review: Basic Ideas of Scientific Sampling, by A. Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (3):211, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986813.

McCormack:1963:BRB

[200] J. McCormack. Book review: Statistical Management of Investory Systems, by Harvey M. Wagner. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):211–212, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986814.

Huitson:1963:BRBd

[201] A. Huitson. Book review: Random Processes, by Murray Rosenblatt. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):212–213, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2986815.

Huitson:1963:BRBe

[202] A. Huitson. Book review: Statistical Methods in Clinical and Preventive Medicine, by Austin Bradford Hill. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (3):213, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986816.

Siskind:1963:BRB

[203] V. Siskind. Book review: Introduction to Statistical Inference, by E. S. Keeping. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (3):213–215, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986817.

Lewis:1963:BRBc

[204] T. W. Lewis. Book review: Wheldon’s Business Statistics and Statistical Methods, by G. L. Thirkettle. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):215–216, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986818.

Anonymous:1963:NAc

[205] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):217– 234 + 236–239, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986819.

Anonymous:1963:BMc

[206] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(3):240–243, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986820.

Ball:1963:CUS

[207] R. J. Ball. The construction and uses of statistical economic models. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):245–256, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987304.

Ehrenberg:1963:SQF

[208] A. S. C. Ehrenberg. Some queries to factor analysts. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):257–262, ???? 1963. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987305.

Solari:1963:DCS

[209] M. E. Solari. The distribution of the chi square test of fit statistic. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):263–267, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987306.

Pridmore:1963:BRB

[210] W. A. Pridmore. Book review: Introductory Descriptive Statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):269–274, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987307.

James:1963:C

[211] A. N. James and Denis H. Ward. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):275–276, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987308.

Farlie:1963:QWT

[212] D. J. Farlie. Quick ways to the top. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4): 277–285, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987309.

Anonymous:1963:Ed

[213] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):287–301, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987310.

Anonymous:1963:FMd

[214] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):302–312, ????

1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987303.

Rimmer:1963:BRB

[215] R. Rimmer. Book review: A Manager’s Guide to Operational Research, by Patrick Rivett; Russell L. Ackoff. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4): 303–304, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987311.

Harrison:1963:BRB

[216] T. K. Harrison. Book review: Numerical Analysis, by Nathaniel Macon. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4): 304–305, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987312.

Huitson:1963:BRBf

[217] A. Huitson. Book review: Contributions to Order Statistics, by A. E. Sarhan; B. G. Greenberg. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):305–306, ???? 1963. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987313.

Huitson:1963:BRBg

[218] A. Huitson. Book review: Handbook of Nonparametric Statistics, by John E. Walsh. Journal of the Royal Statistical Society. Series D (The Statistician), 13 (4):306–307, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987314.

Knowles:1963:BRB

[219] E. A. G. Knowles. Book review: Introduction to Probability and Random Variables, by George P. Wadsworth; Joseph G. Bryan. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):307–309, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987315.

Thomas:1963:BRB

[220] D. A. H. Thomas. Book review: Statistical Analysis-Ideas and Methods, by E. Vernon Lewis; Statistics — An Intuitive Approach, by George H. Weinberg; John A. Schumaker. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):309–311, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987316.

Anonymous:1963:NAd

[221] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4): 313–321, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987317.

Anonymous:1963:BMd

[222] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):322–332, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987319.

Anonymous:1963:AA

[223] Anonymous. Analysis of appointments.

Journal of the Royal Statistical Society. Series D (The Statistician), 13(4): 323–325, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987318.

Anonymous:1963:VI

[224] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 13(4):329–331, ???? 1963. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987302.

Shaw:1964:SCI

[225] D. C. Shaw. Statistics of the cotton industry. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):1–14, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987340.

Holmes:1964:TS

[226] D. Holmes. Training in statistics.

Journal of the Royal Statistical Society. Series D (The Statistician), 14(1): 15–21, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987341.

Low:1964:SAS

[227] J. M. Low. The statistical analysis of stock movements in the man-made fibre industry. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):23–31, ???? 1964. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987342.

Tippett:1964:SII

[228] L. H. C. Tippett. The Shirley Institute, the industry and statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):33–37, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987343.

Norris:1964:MTE

[229] K. P. Norris. Management, technology, and economics at the Shirley Institute. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):39–44, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987344.

Anonymous:1964:FMa

[230] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):45, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987339.

Lindley:1964:FAS

[231] D. V. Lindley, F. W. Warburton, A. S. C. Ehrenberg, J. Jeffers, and D. N. Lawley. Factor analysis: A summary of discussion. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):47–61, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987345.

Hedges:1964:UPA

[232] B. A. A. Hedges. The use of prevaricative analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):63–70, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987346.

Thomas:1964:C

[233] D. A. H. Thomas, William Stevenson, W. A. Pridmore, and E. J. Snell. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):71–72, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987347.

Anonymous:1964:CE

[234] Anonymous. Corrigenda: Education.

Journal of the Royal Statistical Society. Series D (The Statistician), 14 (1):73, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987349.

Anonymous:1964:Ea

[235] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):73, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987348.

Anonymous:1964:RNR

[236] Anonymous. Review notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):74–76, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987350.

Fox:1964:BRB

[237] R. A. Fox. Book review: Introduction to Mathematical Statistics, by Paul G. Hoel. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (1):76–77, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987351.

Huitson:1964:BRBa

[238] A. Huitson. Book review: Geometrical Probability, by M. G. Kendall; P. A. P. Moran. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):77–78, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987352.

Huitson:1964:BRBb

[239] A. Huitson. Book review: Random Wavelets and Cybernetic Systems, by E. A. Robinson. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):78, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987353.

Buckland:1964:BRBa

[240] W. R. Buckland. Book review: Theories of Engineering Experimentation, by H. Schenck. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):78–79, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987354.

Anonymous:1964:NAa

[241] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 14(1): 80–85, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987355.

Anonymous:1964:BMa

[242] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(1):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987356.

Armitage:1964:MS

[243] P. Armitage. Medical statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):87, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987104.

Case:1964:UOS

[244] R. A. M. Case and Joan M. Davies. On the use of official statistics in medical research. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (2):89–119, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987105.

Billewicz:1964:SMI

[245] W. Z. Billewicz. Statistics in medical investigations. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):121–136, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987106.

Spicer:1964:UCM

[246] C. C. Spicer. Use of computers in medical statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):137–145, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987107.

Radhakrishna:1964:DAM

[247] S. Radhakrishna. Discrimination analysis in medicine. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):147–167, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987108.

Anonymous:1964:Eb

[248] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):169–170, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987109.

Thomas:1964:RNR

[249] D. A. H. Thomas. Review notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):171–173, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987110.

Buckland:1964:BRBb

[250] W. R. Buckland. Book review: Statistical Concepts: A Programme for Self-Instruction, by C. McCollough; L. Van Atta. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (2):173–174, ???? 1964. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987111.

Thomas:1964:BRB

[251] D. A. H. Thomas. Book review: The Method of Paired Comparisons, by H. A. David. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (2):175–176, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987112.

Rimmer:1964:BRB

[252] R. Rimmer. Book review: Probability and Statistics for Everyman, by Irving Adler. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):176, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987113.

Wagle:1964:BRB

[253] B. V. Wagle. Book review: Mathematical Programming, by S. Vajda. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2): 176–177, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987114.

Buckland:1964:BRBc

[254] W. R. Buckland. Book review: Statistical Analysis in The Geological Sciences, by R. L. Miller; J. S. Kahn. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2): 177–178, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987115.

Lewis:1964:BRB

[255] T. W. Lewis. Book review: Statistical Measurements for Economics & Administration, by Morris Budin. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):178–179, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987116.

Harrison:1964:BRB

[256] T. K. Harrison. Book review: Bibliography of Basic Texts and Monographs on Statistical Methods, by W. R. Buckland; R. A. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):179–180, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987117.

Huitson:1964:BRBc

[257] A. Huitson. Book review: Renewal Theory, by D. R. Cox. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):180, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987118.

Buckland:1964:BRBd

[258] W. R. Buckland. Book review: Fatigue of Metals, by P. G. Forrest. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2): 180–181, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987119.

Anonymous:1964:NAb

[259] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 14(2): 183–187, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987120.

Anonymous:1964:BMb

[260] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987121.

Anonymous:1964:FMb

[261] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(2):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987103.

Anonymous:1964:F

[262] Anonymous. Foreword. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):189, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987044.

Wollen:1964:SUS

[263] G. H. Wollen, J. E. Outlaw, A. H. J.

Baines, W. J. K. Thomas, K. J. Valentine, E. L. Snowdon, G. Sharp, O. G. Tanner, E. A. G. Shrimpton, and D. D. Filtness. Some uses of statistics within the Ministry of Agriculture, Fisheries and Food. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):191–225, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987045.

Bessell:1964:MLA

[264] J. E. Bessell. Measurement of leadership among farmers. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):227–253, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987046.

Cowling:1964:MSR

[265] Keith Cowling and T. W. Gardner. Milk supply response: An interbreed analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (3):255–266, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987047.

Bessell:1964:TSA

[266] J. E. Bessell. Teaching of statistics to agricultural economists. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):267–273, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987048.

Downham:1964:BRB

[267] J. S. Downham. Book review: British Towns: A Statistical Study of their Social and Economic Differences, by C. A. Moser; Wolf Scott. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):275–276, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987049.

Ilersic:1964:BRB

[268] A. R. Ilersic. Book review: The Use and Abuse of Statistics, by W. J. Reichmann. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):276–277, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987050.

Anonymous:1964:NAc

[269] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 14(3): 278–282, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987051.

Anonymous:1964:BMc

[270] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987052.

Anonymous:1964:FMc

[271] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(3):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987043.

Powell-Jones:1964:IR

[272] D. L. Powell-Jones. Introductory remarks. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (4):283–284, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987018.

Riddell:1964:SRP

[273] G. L. Riddell. “The structure of the Reed Paper Group”. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):285–289, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987019.

Moore:1964:RSS

[274] P. G. Moore. A review of some statistical activities in the Reed paper group. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4): 291–309, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987020.

Draper:1964:SSP

[275] J. Draper. Some statistical problems in research and development. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):311–318, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987021.

Page:1964:SAN

[276] D. J. Page. The statistical assessment of new materials in the manufacture of corrugated board. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):319–329, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987022.

Read:1964:SPA

[277] W. J. Read. Some problems in applying quality control to the packaging industry. Journal of the Royal Statistical Society. Series D (The Statistician), 14 (4):331–334, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987023.

Self:1964:RIM

[278] P. J. Self. The role of industrial market research. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):335–338, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987024.

Blythe:1964:UGC

[279] B. Blythe. The use of the Group’s computer as a computational aid. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):339–340, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987025.

Taylor:1964:C

[280] G. A. R. Taylor and D. V. Lindley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):341–343, ???? 1964. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987026.

Anonymous:1964:ES

[281] Anonymous. Education section. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):345–355, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987027.

Anonymous:1964:AA

[282] Anonymous. Analysis of appointments. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4): 357–359, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987028.

Anonymous:1964:VI

[283] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):360–363, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987016.

Anonymous:1964:BMd

[284] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987029.

Anonymous:1964:FMd

[285] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 14(4):??, ???? 1964. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987017.

Buckland:1965:OKS

[286] Wm. R. Buckland. The organization of knowledge on statistical methods. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):1–5, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987238.

Duckworth:1965:SMM

[287] W. E. Duckworth. Statistical method in metallurgical development. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):7–30, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987239.

Henshaw:1965:SMI

[288] W. C. Henshaw. Statistical methods of inventory control. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):31–42, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987240.

Knight:1965:MVI

[289] G. J. Knight. Measurement of virus infectivity. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):43–57, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987241.

Lewis:1965:SSA

[290] B. D. Lewis. A survey of some of the available methods of analysis of the growth curves of rats in toxicological tests. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):59–66, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987242.

Hays:1965:ICT

[291] S. Hays. International comparisons of the technological factors affecting economic growth. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):67–84, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987243.

Stone:1965:BRB

[292] M. Stone. Book review: Expected Values of Discrete Random Variables and Elementary Statistics, by A. L. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):85, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987244.

Cox:1965:BRB

[293] D. R. Cox. Book review: Statistical Assessment of the Life Characteristic, a Bibliographic Guide, by W. R. Buckland. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):85–86, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987245.

Huitson:1965:BRBa

[294] A. Huitson. Book review: Computer Language — An Autoinstructional Introduction to Fortran, by H. L. Colman; C. Smallwood. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):86–87, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987246.

Huitson:1965:BRBb

[295] A. Huitson. Book review: The Art of Simulation, by K. D. Tocher. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):87–88, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987247.

Kingman:1965:BRB

[296] J. F. C. Kingman. Book review: Inequalities on Distribution Functions, by H. J. Godwin. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):88–89, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987248.

Buckland:1965:BRBa

[297] W. R. Buckland. Book review: Progress in Operations Research. Vol. II, by D. B. Hertz; R. T. Eddison. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):89, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987249.

Oliver:1965:BRB

[298] F. R. Oliver. Book review: British Export Performance, by S. J. Wells. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):90, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987250.

Jeffers:1965:BRB

[299] J. N. R. Jeffers. Book review: Plant Diseases: Epidemics and Control, by J. E. Van Der Plank. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):90–91, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987251.

Ward:1965:BRB

[300] D. H. Ward. Book review: Matrix Algebra for Social Scientists, by Paul Horst. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):91–92, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987252.

Baillie:1965:BRB

[301] D. H. Baillie. Book review: Fundamental Statistics for Business and Economics, by John Neter; William Wasserman. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (1):92–93, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987253.

Anonymous:1965:BRBa

[302] Anonymous. Book review: Basic Statistics: A Primer for the Biomedical Sciences, by Olive Jean Dunn. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):93, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987254.

Henshaw:1965:BRB

[303] W. C. Henshaw. Book review: Game Theory and Related Approaches to Social Behaviour, by Martin Shubik. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):93–96, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987255.

Anonymous:1965:NAa

[304] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 15(1): 97–99, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987256.

Anonymous:1965:BMa

[305] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987257.

Anonymous:1965:FMa

[306] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(1):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987237.

Craddock:1965:SWF

[307] J. M. Craddock. Statistics and weather forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):101–109, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987386.

Glahn:1965:OWF

[308] Harry R. Glahn. Objective weather forecasting by statistical methods. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):111–142, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987387.

Craddock:1965:MAP

[309] J. M. Craddock. A meterological application of principal component analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (2):143–156, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987388.

Freeman:1965:SBG

[310] M. H. Freeman. A statistically based graphical technique of objective forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):157–165, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987389.

Craddock:1965:AMT

[311] J. M. Craddock. The analysis of meteorological time series for use in forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (2):167–190, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987390.

Wagle:1965:RTS

[312] B. V. Wagle. A review of two statistical aids in forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):191–196, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987391.

Fox:1965:BRB

[313] R. A. Fox. Book review: The Statistical Method in Business, by F. A. Ekeblad. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (2):197–198, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987392.

Lindley:1965:BRB

[314] D. V. Lindley. Book review: Principles of Random Walk, by Frank Spitzer. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2): 198–199, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987393.

Huitson:1965:BRBc

[315] A. Huitson. Book review: Introduction to Probability and Statistics from a Bayesian Viewpoint, by D. V. Lindley. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (2):199–200, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987394.

Buckland:1965:BRBb

[316] W. R. Buckland. Book review: Theory and Technique of Variation Research; The Statistical Treatment of Fatigue Experiments, by Leonard G. Johnson. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (2):200, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987395.

Dyson:1965:BRB

[317] R. G. Dyson. Book review: Mathematical Trend Curves — An Aid to Forecasting, by J. V. Gregg; C. H. Hossell; J. T. Richardson. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):201, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987396.

Rimmer:1965:BRB

[318] R. Rimmer. Book review: Short-Term Forecasting, by G. A. Coutie; O. L. Davies; C. H. Hossell; D. W. G. P. Millar; A. J. H. Morrell. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):201, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987397.

Harrison:1965:BRBa

[319] T. K. Harrison. Book review: Cumulative Sum Techniques, by R. H. Woodward; P. L. Goldsmith. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):201–202, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987398.

Anonymous:1965:BRBb

[320] Anonymous. Book review: Fundamental Statistics for Business and Economics, by J. Neter; W. Wasserman; Elements of Statistical Inference, by D. V. Huntsberger. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):202, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987399.

Anonymous:1965:BRBc

[321] Anonymous. Book review: Random Negative Exponential Deviates, by V.

D. Barnett; Fitting of Growth and Allied Curves of the Asymptotic Regression Type by Stevens’ Method, by R. W. Hiorns. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):202, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987400.

Anonymous:1965:BRBd

[322] Anonymous. Book review: Applied Combinatorial Mathematics, by E. F. Beckenbach. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):203, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987401.

Anonymous:1965:BRBe

[323] Anonymous. Book review: Lady Luck:

The Theory of Probability, by Warren Weaver. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):203, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987402.

Anonymous:1965:BRBf

[324] Anonymous. Book review: The Use of Mathematics in Economics, by V. S. Nemchinov. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):203, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987403.

Adler:1965:C

[325] Max K. Adler, P. J. Self, and J. N. R.

Jeffers. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):205–208, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987404.

Anonymous:1965:BMb

[326] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987405.

Anonymous:1965:FMb

[327] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(2):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987385.

Anonymous:1965:SWC

[328] Anonymous. Statistics and weight control problems. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):209–212, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986925.

Ward:1965:WCP

[329] D. H. Ward. The weight control problem — from the point of view of the statistician in industry. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):213–216, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986926.

Gray:1965:WCI

[330] W. Gray. Weight control from the Inspector’s viewpoint. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):217–222, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986927.

Anonymous:1965:SS

[331] Anonymous. Standards and statistics.

Journal of the Royal Statistical Society. Series D (The Statistician), 15(3): 223–227, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986928.

Poppy:1965:LRA

[332] T. G. Poppy. The law relating to automatic weighing of pre-packed commodities. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (3):229–237, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986929.

Lewis:1965:OPM

[333] R. P. Waldo Lewis. Objective pattern matching in long range forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (3):239–245, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986930.

Anonymous:1965:Ea

[334] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):247–250, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986931.

Pluck:1965:BRBa

[335] R. A. Pluck. Book review: Statistics in Theory and Practice, by L. R. Connor; A. J. H. Morrell; Statistics, by A. R. Ilersic. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):251, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986932.

Pluck:1965:BRBb

[336] R. A. Pluck. Book review: Applications of Characteristic Functions, by E. Lukacs; R. G. Laha; Elements on Linear Programming, by R. G. Geary; M. D. McCarthy. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):252, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986933.

Harrison:1965:BRBb

[337] T. K. Harrison. Book review: Elements of Numerical Analysis, by Peter Henrici. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):252–253, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986934.

Barnett:1965:BRB

[338] B. M. Barnett. Book review: Standard Statistical Calculations, by P. G. Moore; D. E. Edwards. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):253, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986935.

Oaten:1965:BRB

[339] M. J. Oaten. Book review: The McGraw-Hill Dictionary of Modern Economics: A Handbook of Terms and Organizations, by Douglas Greenwald. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3): 253–254, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986936.

Lindley:1965:EIP

[340] D. V. Lindley. Erratum: Introduction to probability and statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):254, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986937.

Anonymous:1965:NAb

[341] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 15(3): 255–260, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986938.

Anonymous:1965:IS

[342] Anonymous. The Institute of Statisticians. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (3):261–264 + 266–271, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986939.

Anonymous:1965:BMc

[343] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986940.

Anonymous:1965:FMc

[344] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(3):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986924.

Brech:1965:TF

[345] Ronald Brech. The theory of Fothergale. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (4):273–296, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986749.

Cottrell:1965:SE

[346] H. C. Cottrell. The stock exchange. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4): 297–311, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986750.

Aitchison:1965:SPH

[347] J. Aitchison. Statistical problems in hospital engineering. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):313–318, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986751.

Kasap:1965:SID

[348] H. S. Kasap, Lynda Bates, John Bowles, and A. F. Bissell. The student and the Institute: A discussion. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):319–336, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986752.

Bissell:1965:HHS

[349] A. F. Bissell. Hints for the home student. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (4):337–339, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986753.

Anonymous:1965:DMA

[350] Anonymous. Debate on the motion “That Academic Training is Superior to Training Within Industry”. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):341–343, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986754.

Thomas:1965:C

[351] D. A. H. Thomas and D. V. Lindley.

Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):345–349, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986755.

Huitson:1965:BRBd

[352] A. Huitson. Book review: The Elements of Stochastic Processes with Applications to the Natural Sciences, by N.

T. J. Bailey. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):351, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986756.

Huitson:1965:BRBe

[353] A. Huitson. Book review: Introductory Engineering Statistics, by I. Guttman; S. S. Wilks. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):351–352, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986757.

Elton:1965:BRB

[354] H. D. Elton. Book review: Mathematical Analysis of Observations, by B. M. Shchigolev. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):352–354, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986758.

Doody:1965:BRB

[355] Pamela M. Doody. Book review: Descriptive Statistics, by L. G. Gotkin; L. S. Goldstein. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):354, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986759.

Manwaring:1965:BRB

[356] D. B. Manwaring. Book review: Essentials of Quality Control, by Alan Huitson; Joan Keen. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):354–355, ???? 1965. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986760.

Anonymous:1965:NRE

[357] Anonymous. Notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 15 (4):355–356, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986761.

Anonymous:1965:Eb

[358] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):357–375, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986762.

Anonymous:1965:NAc

[359] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 15(4): 376–378, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986763.

Anonymous:1965:VI

[360] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):379–382, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986747.

Anonymous:1965:BMd

[361] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):??, ???? 1965.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986764.

Anonymous:1965:FMd

[362] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 15(4):??, ???? 1965. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986748.

Anonymous:1966:FMa

[363] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):1, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987072.

Buck:1966:SSA

[364] S. F. Buck and A. J. Wicken. The statistical study of air pollution and smoking. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):3–4, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987073.

Holland:1966:SGD

[365] W. W. Holland. The study of geographic differences in the prevalence of chronic bronchitis. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):5–22, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987074.

Hoinville:1966:ESH

[366] G. W. Hoinville and H. W. Biggs. Establishing smoking habits in retrospect.

Journal of the Royal Statistical Society. Series D (The Statistician), 16(1): 23–43, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987075.

Waller:1966:IUD

[367] R. E. Waller. The interpretation and use of data on air pollution in epidemiological research. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):45–58, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987076.

Buck:1966:PSB

[368] S. F. Buck and A. J. Wicken. A possible source of bias in the calculation of mortality rates for various smoking habit classes. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):59–68, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987077.

Craddock:1966:SCL

[369] J. M. Craddock. A simple computer language for the statistician. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):69–85, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987078.

Craddock:1966:TSC

[370] J. M. Craddock. Testing the significance of a 3 × 3 contingency table. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1): 87–94, ???? 1966. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987079.

Baillie:1966:BRB

[371] D. H. Baillie. Book review: Statistics; an Introduction for Students in the Behavioural Sciences, by Dorothy D. Adkins; Statistical Methods: A Problem-Solving Approach, by V. Gourevitch. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1): 95–96, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987080.

Doody:1966:BRB

[372] Pamela M. Doody. Book review: Definitions and Formulae for Students: Statistics, by R. W. Hiorns. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):96, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987081.

Manwaring:1966:BRBa

[373] D. B. Manwaring. Book review: The Statistical Analysis of Experimental Data, by John Mandel; A Statistical Study of Physical Classroom Experiments, by John Mandel. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):97–98, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987082.

Buckland:1966:BRB

[374] W. R. Buckland. Book review: Random Processes and the Growth of Firms, by J. Steindl. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):98–99, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987083.

Anonymous:1966:Ea

[375] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):101–113, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987084.

Anonymous:1966:NAa

[376] Anonymous. News and announcements.

Journal of the Royal Statistical Society. Series D (The Statistician), 16(1): 114–117, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987085.

Anonymous:1966:BMa

[377] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(1):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987086.

Taylor:1966:RQW

[378] S. M. Taylor. Recent quantitative work on British bird populations. a review. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2): 119–170, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986734.

Boyd:1966:AEW

[379] Hugh Boyd. The assessment of the effects of weather upon the breeding success of geese nesting in the Arctic. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2): 171–181, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986735.

Murton:1966:SEE

[380] R. K. Murton. A statistical evaluation of the effect of wood-pigeon shooting as evidenced by the recoveries of ringed birds. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (2):183–202, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986736.

Calder:1966:SAS

[381] A. B. Calder. Some applications of statistical procedure in analytical chemistry. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (2):203–216, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986737.

Harrison:1966:BRB

[382] T. K. Harrison. Book review: Principles of Statistics, by M. G. Bulmer. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):217, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986738.

Rimmer:1966:BRBa

[383] R. Rimmer. Book review: Concepts of Statistical Inference, by William C. Guenther. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):217, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986739.

Rimmer:1966:BRBb

[384] R. Rimmer. Book review: An Introduction to Quantitative Business Analysis, by Ira Horwitz. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):218, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986740.

Elton:1966:BRBa

[385] H. D. Elton. Book review: A Guide to FORTRAN IV Programming, by D. D. McCracken. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):218–219, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986741.

Huitson:1966:BRBa

[386] A. Huitson. Book review: Queues and Inventories, by N. U. Prabhu. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2): 219–220, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986742.

Huitson:1966:BRBb

[387] A. Huitson. Book review: Handbook of Nonparametric Statistics, Vol II, by John E. Walsh. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):220, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986743.

Anonymous:1966:BRBa

[388] Anonymous. Book review: Index Numbers, Theory and Applications, by W. R. Crowe. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (2):220–221, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986744.

Anonymous:1966:NAb

[389] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2): 223–226, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986745.

Anonymous:1966:BMb

[390] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986746.

Anonymous:1966:FMb

[391] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(2):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986733.

Goldsmith:1966:TPS

[392] P. L. Goldsmith. The theoretical performance of a semi-continuous blending system. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):227–251, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986871.

Wagle:1966:STS

[393] B. Wagle. Some techniques of shortterm sales forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):253–273, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986872.

Liddell:1966:HSS

[394] F. D. K. Liddell. The health of senior staff. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (3):275–296, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986873.

Wood:1966:SCT

[395] P. D. P. Wood. A suggested computer technique for the analysis of nonorthogonal data in a multi-way table. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3): 297–303, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986874.

Fox:1966:BRBa

[396] R. A. Fox. Book review: Survey Sampling, by Leslie Kish. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):305–306, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986875.

Fox:1966:BRBb

[397] R. A. Fox. Book review: Econometrics, by Gerhard Tintner. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):306, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986876.

Fox:1966:BRBc

[398] R. A. Fox. Book review: Glossary of Terms in Official Statistics: English-French, French-English, by J. W. Nixon. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (3):306, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986877.

Huitson:1966:BRBc

[399] A. Huitson. Book review: The Statistical Analysis of Series of Events, by D. R. Cox; P. A. W. Lewis. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):306–307, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986878.

Manwaring:1966:BRBb

[400] D. B. Manwaring. Book review: Introduction to Statistical Ideas for Social Scientists, by Graham Kalton. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):307–308, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986879.

Ford:1966:BRB

[401] S. J. Ford. Book review: Studies in Subjective Probability, by H. E. Kyburg,; H. E. Smokler. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):308–309, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986880.

Kendall:1966:BRB

[402] D. G. Kendall. Book review: Bibliography on Time Series and Stochastic Processes, by Herman O. A. Wold. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3): 309–311, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986881.

Anonymous:1966:BRBb

[403] Anonymous. Book review: Statistics.

An Introduction to Quantitative Economic Research, by Daniel B. Suits. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (3):311, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986882.

Elton:1966:BRBb

[404] H. D. Elton. Book review: Matrix Algebra for the Biological Sciences (Including Applications in Statistics), by S. R. Searle. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):312, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986883.

Anonymous:1966:NRE

[405] Anonymous. Notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (3):312–313, ???? 1966. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986884.

Clough:1966:C

[406] W. A. Clough and D. V. Lindley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):314–316, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986885.

Anonymous:1966:O

[407] Anonymous. Obituaries. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):317–318, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986886.

Anonymous:1966:ESA

[408] Anonymous. Education and student affairs. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (3):319–321, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986887.

Anonymous:1966:NAc

[409] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3): 322–324, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986888.

Anonymous:1966:RCY

[410] Anonymous. Report of Council for the year 1965 to the Seventeenth Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):325–334, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986889.

Anonymous:1966:ISM

[411] Anonymous. The Institute of Statisticians minutes of the 17th Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):335–336, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986890.

Anonymous:1966:BMc

[412] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986891.

Anonymous:1966:FMc

[413] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(3):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986870.

Anonymous:1966:Eb

[414] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):337–338, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987182.

Clementson:1966:SAI

[415] A. T. Clementson. Simulation applied to industry. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):339–350, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987183.

Myers:1966:FCP

[416] C. L. Myers. Forecasting Child populations. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (4):351–369, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987184.

Davis:1966:JBJ

[417] E. J. Davis. Juke Box jury as a predictive instrument. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):371–379, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987185.

Moore:1966:SC

[418] P. G. Moore. A survey of consignments. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (4):381–386, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987186.

Sargan:1966:BRB

[419] J. D. Sargan. Book review: On the Theory and Measurement of Technological Change, by Murray Brown. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):387, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987187.

Anonymous:1966:BRBc

[420] Anonymous. Book review: Statistics for Business, by Leonard J. Holman. Journal of the Royal Statistical Society. Series D (The Statistician), 16 (4):388, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987188.

Manwaring:1966:BRBc

[421] D. B. Manwaring. Book review: The Mathematical Theory of Reliability, by R. E. Barlow; F. Proschan; L. C. Hunter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4): 388–389, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987189.

Elton:1966:BRBc

[422] H. D. Elton. Book review: Elementary Statistics, by P. G. Hoel. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):389–390, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987190.

Armitage:1966:BRB

[423] P. Armitage. Book review: Biological Statistics: An Introduction, by S. C. Pearce. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):390, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987191.

Newton:1966:C

[424] R. G. Newton and D. C. Blackley. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):391–392, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987192.

Anonymous:1966:Ec

[425] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):393–400, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987193.

Anonymous:1966:EER

[426] Anonymous. 1966 examinations examiners’ reports. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):401–419, ???? 1966. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987194.

Anonymous:1966:NAd

[427] Anonymous. News and announcements. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4): 421–427, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987195.

Anonymous:1966:VI

[428] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):429–432, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

                                        URL http://www.jstor.org/stable/
                                        2987180.

Anonymous:1966:BMd

[429] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987196.

Anonymous:1966:FMd

[430] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 16(4):??, ???? 1966. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987181.

Sneath:1967:SSP

[431] P. H. A. Sneath. Some statistical problems in numerical taxonomy. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):1–12, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987198.

Gower:1967:MAM

[432] J. C. Gower. Multivariate analysis and multidimensional geometry. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):13–28, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987199.

Jeffers:1967:SVT

[433] J. N. R. Jeffers. The study of variation in taxonomic research. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):29–43, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987200.

Needham:1967:ACL

[434] R. M. Needham. Automatic classification in linguistics. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):45–54, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987201.

Pridmore:1967:OSD

[435] W. A. Pridmore. The organization of a statistical department in manufacturing industry. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):55–70, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987202.

Anonymous:1967:I

[436] Anonymous. [introduction]. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):71, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987203.

Franklin:1967:SPM

[437] M. Franklin. Survey on pocket money. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1): 73–80, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987204.

Moore:1967:BRB

[438] P. G. Moore. Book review: Linear Statistical Inference and its Applications, by C. Radhakrishna Rao. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):81–82, ???? 1967.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987205.

Blackmore:1967:BRB

[439] R. W. Blackmore. Book review: Mathematics and Statistics for Scientists and Engineers, by P. MacDonald; Mathematics and Statistics for Technologists, by H. G. Cuming; C. J. Anson. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (1):82–83, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987206.

Campbell:1967:BRB

[440] R. C. Campbell. Book review: Logic of Statistical Inference, by I. Hacking. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1): 83–85, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987207.

Garza:1967:BRB

[441] Tomas Garza. Book review: Green’s Function Methods in Probability Theory, by J. Keilson. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):85, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987208.

Huitson:1967:BRBa

[442] A. Huitson. Book review: Sequential Methods in Statistics, by G. B. Wetherill. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (1):85–86, ???? 1967. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987209.

Huitson:1967:BRBb

[443] A. Huitson. Book review: The Theory of Stochastic Processes, by D. R. Cox; H. D. Miller. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):86–87, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987210.

Foister:1967:BRB

[444] A. F. M. Foister. Book review: Statistical Aids to Railway Operation, by Jagjit Singh. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (1):87–88, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987211.

Anonymous:1967:NREa

[445] Anonymous. Notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (1):88–92, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987212.

Davies:1967:C

[446] P. M. Davies. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):93, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987213.

Anonymous:1967:EER

[447] Anonymous. 1966 examinations examiners’ reports. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):95–106, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987214.

Anonymous:1967:Ea

[448] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):107–112, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987215.

Anonymous:1967:Sa

[449] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):113, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987216.

Anonymous:1967:Ma

[450] Anonymous. Members. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):114–117, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987217.

Anonymous:1967:Ga

[451] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):118–120, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987218.

Craddock:1967:EAM

[452] J. M. Craddock. Erratum: The analysis of meteorological time series for use in forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):121, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987219.

Anonymous:1967:BMa

[453] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987220.

Anonymous:1967:FMa

[454] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(1):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987197.

Anonymous:1967:Eb

[455] Anonymous. Editorial. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):123, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986893.

Cranston:1967:ISB

[456] R. W. Cranston. Introduction to statistics at the BP research centre. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2): 125–126, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986894.

OCarroll:1967:PEM

[457] F. M. O’Carroll. Precision evaluation and monitoring. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):127–140, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986895.

Jenkins:1967:MMA

[458] G. I. Jenkins. Multivariate methods applied to product testing and specifications. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (2):141–155, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986896.

Shelton:1967:SEP

[459] R. J. Shelton. Stochastic experimentation principles and practice. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):157–181, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986897.

Williams:1967:EET

[460] J. D. Williams. Evaluation in education:

Ten dilemmas of design. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):183–191, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986898.

Brown:1967:GP

[461] B. C. Brown and O. Nankivell. The Genesis of the plan. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):193–196, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986899.

Ward:1967:BRB

[462] D. H. Ward. Book review: An Introduction to Probability Theory and its Application, Vol. II, by W. Feller. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2): 197–198, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986900.

Huddie:1967:BRB

[463] R. F. Huddie. Book review: The Analysis of Variance, by A. Huitson. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):198, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986901.

Elton:1967:BRB

[464] H. D. Elton. Book review: Modern Elementary Statistics, by John F. Freund; Elements of Mathematical Statistics, by J. F. Ractliffe. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):198–199, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986902.

Treasure:1967:BRB

[465] J. Treasure. Book review: Marketing and Market Assessment, by J. L. Sewell. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2): 199–200, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986903.

Anonymous:1967:RCY

[466] Anonymous. Report of Council for the Year 1966 to the Eighteenth Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):201–211, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986904.

Anonymous:1967:Ec

[467] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):212–223, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986905.

Anonymous:1967:Mb

[468] Anonymous. Members. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):224–225, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986906.

Anonymous:1967:Sb

[469] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):226–229, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986907.

Anonymous:1967:Gb

[470] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):230–232, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986908.

Anonymous:1967:BMb

[471] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986909.

Anonymous:1967:FMb

[472] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(2):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986892.

Bowyer:1967:I

[473] B. F. Bowyer. [introduction]. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):233–236, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986716.

VanDobbenDeBruyn:1967:ITS

[474] C. S. Van Dobben De Bruyn. The interplay of tracking signals and adaptive predictors. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):237–246, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986717.

Burman:1967:ASA

[475] J. P. Burman. Assessment of a seasonal adjustment procedure by spectral analysis. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (3):247–256, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986718.

Craddock:1967:EAP

[476] J. M. Craddock. An experiment in the analysis and prediction of time series. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (3):257–268, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986719.

Davies:1967:SFC

[477] Roger Davies and Alan Huitson. A sales forecasting comparison. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):269–278, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986720.

Hughes:1967:UCF

[478] P. A. B. Hughes and N. B. Morgan.

The use of computers in forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3): 279–299, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986721.

Smith:1967:CSM

[479] T. M. F. Smith. A comparison of some models for predicting time series subject to seasonal variation. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):301–305, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986722.

Wood:1967:CME

[480] P. D. P. Wood. A computational method of estimating variance components from non-orthogonal data in a mixed model. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3): 307–314, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986723.

Baillie:1967:BRB

[481] D. H. Baillie. Book review: Problems for Computer Solution, by Fred Gruenberger; George Jaffray. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):315, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986724.

Buckland:1967:BRBa

[482] W. R. Buckland. Book review: Econometric Models and Methods, by C. F. Christ. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (3):315–316, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986725.

Harrison:1967:BRB

[483] T. K. Harrison. Book review: Handbook of Tables for Probability and Statistics, by William H. Beyer. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):316–317, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986726.

Grodman:1967:BRB

[484] L. G. Grodman. Book review: Programming the IBM System/360, by Staff of Computer Usage Company; Ascher Opler. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):317, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986727.

Ogborn:1967:BRB

[485] M. G. Ogborn. Book review: An Introduction to English Historical Demography, by E. A. Wrigley. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):318–319, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986728.

Anonymous:1967:Ed

[486] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):320–323, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986729.

Anonymous:1967:Sc

[487] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):324–328, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986730.

Anonymous:1967:Gc

[488] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):329–333, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986731.

Anonymous:1967:BMc

[489] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986732.

Anonymous:1967:FMc

[490] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(3):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986715.

Brown:1967:ETS

[491] R. V. Brown. Evaluation of total survey error. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):335–356, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987089.

Herne:1967:HCR

[492] H. Herne. How to Cook relationships. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4): 357–370, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987090.

Hay:1967:NSD

[493] W. A. Hay. Non-standard designs. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4): 371–384, ???? 1967. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987091.

Bissell:1967:SAA

[494] A. F. Bissell. “Subjective aspects of applied statistics”. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):385–400, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987092.

Moroney:1967:EI

[495] M. J. Moroney. Ethics and the Institute. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (4):401–405, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987093.

Staples:1967:EEM

[496] G. E. Staples. Efficiency in environmental measurement. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):407–420, ???? 1967. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987094.

Buckland:1967:BRBb

[497] W. R. Buckland. Book review: Elements of Nonparametric Statistics, by G. E. Noether. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):421, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987095.

Anonymous:1967:BRB

[498] Anonymous. Book review: A First Course in Statistics, by Robert Loveday. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (4):422, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987096.

Barton:1967:BRB

[499] D. E. Barton. Book review: Bibliography of Statistical Literature 1940– 1949, by M. G. Kendall; Alison G. Doig. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4): 422–423, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987097.

Anonymous:1967:NREb

[500] Anonymous. Notes by the Review Editor. Journal of the Royal Statistical Society. Series D (The Statistician), 17 (4):423–424, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987098.

Anonymous:1967:Ee

[501] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):425–433, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987099.

Anonymous:1967:Sd

[502] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):434–439, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987100.

Anonymous:1967:Gd

[503] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):440–444, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987101.

Anonymous:1967:VI

[504] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):445–448, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987087.

Anonymous:1967:BMd

[505] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987102.

Anonymous:1967:FMd

[506] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 17(4):??, ???? 1967. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987088.

Pridmore:1968:OST

[507] W. A. Pridmore. The organization of a sensory testing panel. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):1–10, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987160.

Hamilton:1968:SNR

[508] Max Hamilton. Some notes on rating scales. Journal of the Royal Statistical Society. Series D (The Statistician), 18 (1):11–17, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987161.

Taylor:1968:ISP

[509] J. Taylor. The influence of statistical principles on the design and conduct of sensory tests of foodstuffs. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):19–23, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987162.

Morris:1968:FT

[510] M. Morris. A family topic. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):25–30, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987163.

Williams:1968:CSA

[511] W. T. Williams and G. N. Lance. Choice of strategy in the analysis of complex data. Journal of the Royal Statistical Society. Series D (The Statistician), 18 (1):31–43, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987164.

Anonymous:1968:EER

[512] Anonymous. 1967 examinations examiners’ reports. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):45–67, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987165.

Blackmore:1968:BRB

[513] R. Blackmore. Book review: Statistics, an Introduction, by Rickmers. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):69, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987166.

Rimmer:1968:BRB

[514] R. Rimmer. Book review: Applied Regression Analysis, by N. R. Draper; H. Smith. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):69–70, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987167.

Cooper:1968:BRB

[515] B. E. Cooper. Book review: Exercises in Probability and Statistics for Mathematics Undergraduates, by N. A. Rahman. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):70, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987168.

Parker:1968:BRB

[516] J. B. Parker. Book review: Stationary and Related Stochastic Processes, by H. Cramer; M. R. Leadbetter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):70–71, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987169.

Buckland:1968:BRBa

[517] W. R. Buckland. Book review: J. Neyman: Selection of Early Statistical Papers, by J. Neyman. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):71–72, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987170.

Huitson:1968:BRB

[518] A. Huitson. Book review: Computer Programming and Related Mathematics, by R. V. Andree. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):72–73, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987171.

Fox:1968:BRBa

[519] R. A. Fox. Book review: Statistical Methods and Formulae, by C. G. Lambe. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):73, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987172.

Fox:1968:BRBb

[520] R. A. Fox. Book review: An Introduction to Management Statistics, by Programme Division of International Tutor Machines Ltd. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):73–74, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884

                                        (electronic). URL http://www.jstor.
                                        org/stable/2987173.

OCarroll:1968:C

[521] F. M. O’Carroll and J. M. Craddock.

Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):75, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987174.

Anonymous:1968:Ea

[522] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):77–79, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987175.

Anonymous:1968:Sa

[523] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):80–88, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987176.

Anonymous:1968:M

[524] Anonymous. Members. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):89–90, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987177.

Anonymous:1968:Ga

[525] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):91–92, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987178.

Anonymous:1968:BMa

[526] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987179.

Anonymous:1968:FMa

[527] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(1):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987159.

Goldstein:1968:LSM

[528] H. Goldstein. Longitudinal studies and the measurement of change. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):93–117, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986775.

Skurnik:1968:DRE

[529] Larry S. Skurnik and Desmond L. Nuttal. Describing the reliability of examinations. Journal of the Royal Statistical Society. Series D (The Statistician), 18 (2):119–128, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986776.

Robertson:1968:SPC

[530] T. S. Robertson. Some practical considerations in the conduct of educational research in the United Kingdom. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2): 129–140, ???? 1968. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986777.

Wagle:1968:PST

[531] B. Wagle, J. Q. G. H. Rappoport, and V. A. Downes. A program for shortterm sales forecasting. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):141–147, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986778.

Watts:1968:FTB

[532] G. Watts. Flavour tests on beer.

Journal of the Royal Statistical Society. Series D (The Statistician), 18(2): 149–156, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986779.

Sparrow:1968:DS

[533] A. E. Sparrow. Drinking survey. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):157–161, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986780.

Anonymous:1968:BRB

[534] Anonymous. Book review: The Advanced Theory of Statistics, by M. G. Kendall; A. Stuart. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):163–164, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986781.

Buckland:1968:BRBb

[535] W. R. Buckland. Book review: S. S. Wilks: Collected Papers, by S. S. Wilks; T. W. Anderson. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):164–165, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986782.

Elton:1968:BRBa

[536] H. D. Elton. Book review: Elementary Sampling Theory, by Taro Yamane. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2): 165–166, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2986783.

Buckland:1968:BRBc

[537] W. R. Buckland. Book review: Stochastic Models for Social Processes, by D. J. Bartholomew. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):166–168, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986784.

Anonymous:1968:Eb

[538] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):169–181, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986785.

Anonymous:1968:Sb

[539] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):183–188, ???? 1968.

CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986786.

Anonymous:1968:IM

[540] Anonymous. Institute and members.

Journal of the Royal Statistical Society. Series D (The Statistician), 18(2): 189–200, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986787.

Anonymous:1968:Gb

[541] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):201–206, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986788.

Anonymous:1968:BMb

[542] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986789.

Anonymous:1968:FMb

[543] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(2):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986774.

Page:1968:IT

[544] D. J. Page. Innovation in teaching.

Journal of the Royal Statistical Society. Series D (The Statistician), 18 (3):207, ???? 1968. CODEN ????

ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987293.

Kitchen:1968:TMS

[545] D. Kitchen. Teaching management statistics by programmed instruction. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3): 209–226, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987294.

Davies:1968:PCS

[546] Hilda M. Davies. Practical classes in statistics. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):227–236, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987295.

Bissell:1968:LTS

[547] A. F. Bissell and T. F. J. Hobson. Learning and teaching statistics by correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 18 (3):237–244, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987296.

Taylor:1968:TST

[548] S. M. Taylor. Teaching statistics in a technical College. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):245–252, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987297.

Anonymous:1968:ER

[549] Anonymous. Examiners’ reports. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):253–301, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987298.

Anonymous:1968:Ec

[550] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):303–308, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987299.

Anonymous:1968:Gc

[551] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):309–311, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987300.

Anonymous:1968:BMc

[552] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):312, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987301.

Anonymous:1968:FMc

[553] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(3):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987292.

Lindley:1968:DM

[554] D. V. Lindley. Decision making. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):313–326, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987135.

Neale:1968:USL

[555] A. B. Neale. The use of statistics in local government. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):327–344, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987136.

Fry:1968:RIU

[556] R. E. Fry and F. M. Martin. The research and intelligence unit of the Greater London Council. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):345–354, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987137.

Goldsmith:1968:CLD

[557] P. L. Goldsmith. The construction of life distributions from garment wear-trials. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4): 355–375, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987138.

Stewart:1968:DCF

[558] Duncan H. Stewart. Data capture in forestry research. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):377–411, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987139.

Westlake:1968:BRBa

[559] A. J. Westlake. Book review: The Design and Analysis of Scientific Experiments, by K. C. Peng. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):413–414, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987140.

Buckland:1968:BRBd

[560] W. R. Buckland. Book review: Statistical Methods, by G. W. Snedecor; W. G. Cochran. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):414–415, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987141.

Elton:1968:BRBb

[561] H. D. Elton. Book review: Computerized Economic Analysis, by Lawrence Salzman. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):415–417, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987142.

Westlake:1968:BRBb

[562] A. J. Westlake. Book review: Introduction to Probability and Statistics, by W. Mendenhall. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):417, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2987143.

Buckland:1968:BRBe

[563] W. R. Buckland. Book review: World Population Conference: Volume I. Summary Report; World Population Conference: Volume II. Fertility, Family Planning and Mortality; World Population Conference: Volume III. Projection, Population Trends; World Population Conference: Volume IV. Migration, Urbanization and Economic Development (to appear). Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):418–420, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987144.

Holgate:1968:BRB

[564] P. Holgate. Book review: Combinatorial Methods in the Theory of Stochastic Processes, by L. Takacs. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):420–421, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987145.

Silver:1968:BRB

[565] W. E. Silver. Book review: Operational Research, by M. S. Makower; E. Williamson. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):421–422, ???? 1968. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987146.

Anonymous:1968:Sc

[566] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):423–435, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987147.

Anonymous:1968:VI

[567] Anonymous. Volume information. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):437–440, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987133.

Anonymous:1968:BMd

[568] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987148.

Anonymous:1968:FMd

[569] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 18(4):??, ???? 1968. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987134.

McKennell:1969:MPS

[570] A. C. McKennell. Methodological problems in a survey of aircraft noise annoyance. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):1–29, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986700.

Holmes:1969:CSS

[571] C. Holmes. Construction and stratification of a sampling frame of primary sampling units. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):31–47, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986701.

Clunies-Ross:1969:PAT

[572] C. W. Clunies-Ross. Profiles and association: Two exploratory techniques in the multivariate analysis of surveys. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1): 49–60, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986702.

Jenney:1969:TSQ

[573] B. W. Jenney and D. W. Newton. The teaching of statistics for quality control. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (1):61–72, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986703.

Greenhalgh:1969:C

[574] C. Greenhalgh. Correspondence. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):73, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2986704.

Gammon:1969:BRBa

[575] J. Gammon. Book review: Handbook of Methods of Applied Statistics, by I. M. Chakravarti; R. G. Laha; J. Roy. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):75, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986705.

Gammon:1969:BRBb

[576] J. Gammon. Book review: Mathematical Model Building in Economics and Industry, by C. E. I. R. Ltd. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):76, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986706.

Buckland:1969:BRBa

[577] W. R. Buckland. Book review: Introduction to Stochastic Processes in Biostatistics, by C. L. Chiang. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):77, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986707.

Westlake:1969:BRBa

[578] A. J. Westlake. Book review: Solving Problems in Statistics I, by J. Aitcheson. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (1):77–78, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2986708.

Anonymous:1969:BRBa

[579] Anonymous. Book review: The Organization of Knowledge. Industrial Scheduling Abstracts (1950–1966), by S. Eilon; J. R. King; Inventory Control Abstracts (1953–1965), by S. Eilon; W. Lampkin; Critical Path Analysis of Practice, by V. G. Thornley. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):78–80, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884

(electronic). URL http://www.jstor.
org/stable/2986709.

Anonymous:1969:Ea

[580] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):81–85, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986710.

Anonymous:1969:Sa

[581] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):87–97, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986711.

Anonymous:1969:M

[582] Anonymous. Members. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):98, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986712.

Anonymous:1969:G

[583] Anonymous. General. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):99–104, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986713.

Anonymous:1969:BMa

[584] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):??, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2986714.

Anonymous:1969:FMa

[585] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(1):??, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2986699.

Fessey:1969:SDE

[586] M. C. Fessey. Some developments in economic statistics since 1934. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):105–123, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic).

URL http://www.jstor.org/stable/
2987269.

Cooper:1969:SCP

[587] B. E. Cooper. Statistical computing — past, present, and the future. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2): 125–141, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987270.

Newton:1969:EAA

[588] R. G. Newton and D. J. Spurrell. Element analysis — the anatomy of multiple regression. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):143–151, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987271.

Hobson:1969:RAC

[589] T. F. J. Hobson. Regression analysis with categorical regressor variables. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2): 153–161, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987272.

Wood:1969:PVP

[590] P. D. P. Wood. The predictive value of partial sums. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):163–170, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987273.

Buckland:1969:BRBb

[591] W. R. Buckland. Book review: Intermediate Economic Statistics, by Karl A. Fox. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):171–172, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987274.

Elton:1969:BRB

[592] H. D. Elton. Book review: Problems of Stocks and Storage, by A. J. H. Morrell. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):172–173, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987275.

Blight:1969:BRB

[593] B. J. N. Blight. Book review: Advanced Seminar on Spectral Analysis of Time Series, by Bernard Harris. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):173–175, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987276.

Anonymous:1969:BRBb

[594] Anonymous. Book review: Bayesian Decision Problems and Markov Chains, by J. J. Martin. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):175–176, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987277.

Lindley:1969:BRB

[595] D. V. Lindley. Book review: An Introduction to Bayesian Statistical Decision Processes, by Bruce W. Morgan. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2): 177–178, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987278.

Barton:1969:BRB

[596] D. E. Barton. Book review: Bibliography of Statistical Literature, by M. G. Kendall; Alison Doig. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):178–180, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987279.

Anonymous:1969:BRBc

[597] Anonymous. Book review: Decision Analysis: Introductory Lectures on Choices Under Uncertainty, by Howard Raiffa. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):180–181, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987280.

Lanch:1969:BRB

[598] C. Lanch. Book review: Can You Manage Statistics?, by G. E. G. Campling; The Machinery Publishing Co.

Ltd. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):181, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987281.

Westlake:1969:BRBb

[599] A. J. Westlake. Book review: Introduction to Linear Models and the Design and Analysis of Experiments, by W. Mendenhall. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):181–182, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987282.

Fox:1969:BRBa

[600] R. A. Fox. Book review: Elementary Statistical Methods, by G. Barrie Wetherill. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):182–183, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

                                        tronic). URL http://www.jstor.org/
                                        stable/2987283.

Fox:1969:BRBb

[601] R. A. Fox. Book review: The Construction and Use of Economic Models, by A. R. Bergstrom. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):183–184, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987284.

Cooper:1969:BRB

[602] B. E. Cooper. Book review: Multivariate Statistical Methods, by Donald F. Morrison. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):184–185, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987285.

Westlak:1969:BRB

[603] A. J. Westlak. Book review: Statistics for the Social Scientist. Vol. 1 Introducing Statistics; Statistics for the Social Scientist: Vol. 2 Applied Statistics, by K. A. Yeomans. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):185–186, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987286.

Silver:1969:BRB

[604] W. E. Silver. Book review: The Spell of Mathematics, by W. J. Reichmann. Journal of the Royal Statistical Society. Series D (The Statistician), 19 (2):186, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (elec-

tronic). URL http://www.jstor.org/
stable/2987287.

Anonymous:1969:RCY

[605] Anonymous. Report of Council for the year 1969 to the Twentieth Annual General Meeting. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):187–197, ???? 1969. CO-DEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987288.

Anonymous:1969:Eb

[606] Anonymous. Education. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):198–204, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

www.jstor.org/stable/2987289.

Anonymous:1969:Sb

[607] Anonymous. Status. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):205–214, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987290.

Anonymous:1969:BMb

[608] Anonymous. Back matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):??, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987291.

Anonymous:1969:FMb

[609] Anonymous. Front matter. Journal of the Royal Statistical Society. Series D (The Statistician), 19(2):??, ???? 1969. CODEN ???? ISSN 0039-0526 (print), 1467-9884 (electronic). URL http://

                                        www.jstor.org/stable/2987268.